### v1

# Temporal Causal Discovery in Dynamic Bayesian Networks Using Federated Learning

Abstract
Traditionally, learning the structure of a Dynamic Bayesian Network has been centralized, with all data pooled in one location. However, in real-world scenarios, data are often dispersed among multiple parties (e.g., companies, devices) that aim to collaboratively learn a Dynamic Bayesian Network while preserving their data privacy and security. In this study, we introduce a federated learning approach for estimating the structure of a Dynamic Bayesian Network from data distributed horizontally across different parties. We propose a distributed structure learning method that leverages continuous optimization so that only model parameters are exchanged during optimization. Experimental results on synthetic and real datasets reveal that our method outperforms other state-of-the-art techniques, particularly when there are many clients with limited individual sample sizes.
1 Introduction
The learning of Dynamic Bayesian Network (DBN) structures is a crucial technique in causal learning and representation learning. More importantly, it has been applied in various real-world applications, such as modeling gene expression [21], measuring network security [8], studying manufacturing process parameters [47], and recognizing interaction activities [54]. Traditionally, the learning of DBNs has been conducted at a centralized location where all data is aggregated. However, with the rapid advancements in technology and the Internet of Things (IoT) [27] devices, data collection has become significantly more accessible. Consequently, in real-world scenarios, data is typically owned and dispersed among multiple entities, including mobile devices, individuals, companies, and healthcare institutions.
In many cases, these entities, referred to as clients, may lack an adequate number of samples to independently construct a meaningful DBN. They might seek to collaboratively gain comprehensive insights into the DBN structure without disclosing their individual data due to privacy and security concerns. For instance, consider several hospitals aiming to work together to develop a DBN that uncovers the conditional independence relationships among their variables of interest, such as various medical conditions. Sharing patients’ records is clearly prohibited by privacy regulations [2]. Consequently, the sensitive nature of the data renders the centralized aggregation of data from different clients for DBN learning infeasible.
A robust privacy-preserving learning strategy is therefore essential. In recent years, federated learning (FL) [31] has gained significant attention, enabling numerous clients to collaboratively train a machine learning model through a coordinated and privacy-preserving approach. In this framework, each client is not required to share its raw data; instead, they only provide the minimal necessary information, such as model parameters and gradient updates, pertinent to the specific learning task. This approach has been applied across various research domains, including the Internet of Things [20] and recommender systems [7]. Readers may refer to these review papers for more comprehensive insights [23, 51].
Traditionally, score-based learning of DBNs relies on discrete optimization, making it incompatible with common continuous optimization approaches used in federated learning methods. Fortunately, recent work by Zheng et al. [55] and Pamfil et al. [36] have made continuous optimization strategies possible for DBN learning by providing algebraic characterizations of acyclicity. Despite this advancement, directly applying popular federated optimization techniques (FedAvg [31], Ditto [22]) to these continuous formulations remains challenging due to the acyclicity constraint inherent in DBN structure learning.
Contributions.
In this work, we present a federated learning approach for estimating the structure of Dynamic Bayesian Network from observational data (with temporal information) that is horizontally partitioned across different agents. Our contributions are threefold.
  • We propose a Federated DBN Learning method 2Dbn based on continuous optimization using the Alternating Direction Method of Multipliers (ADMM). 2Dbn can simultaneously learn contemporaneous and time-lagged causal structure without any implicit assumptions on the underlying graph topology.
  • Using extensive simulations and two real-world, high-dimensional datasets, we demonstrate that 2Dbn is scalable and effectively identifies the true underlying DBN graph.
  • To our knowledge, this is the first study to explore learning DBN within the FL framework.
Our full code is available on GitHub: https://github.com/PeChen123/2DBN_learning.
Literature Review.
Dynamic Bayesian Network are probabilistic models that extend static Bayesian Networks to represent multi-stage processes by modeling sequences of variables over different stages. Learning DBN involves two key tasks: structure learning (identifying the network topology) and parameter learning (estimating the conditional probabilities). Friedman et al. [12] laid the foundation for this field by developing methods to learn the structure of dynamic probabilistic networks, effectively extending Bayesian Networks to capture temporal dynamics in sequential data. Building on this work, Murphy [32] adapted static Bayesian Network learning techniques for temporal domains, employing methods such as the Expectation-Maximization algorithm for parameter estimation. Advancements in structure learning from time series data were further propelled by Pamfil et al. [36], who developed an extension of Bayesian Network inference using continuous optimization methods. Their approach enabled efficient and scalable learning of DBN structures by formulating the problem as a continuous optimization task with acyclicity constraints tailored for temporal graphs. To address non-linear dependencies in time series data, Tank et al. [48] introduced neural Granger causality methods, which leverage neural networks within the DBN framework to capture complex non-linear interactions. In the field of bioinformatics, Yu et al. [52] enhanced Bayesian Network inference techniques to generate causal networks from observational biological data. Their work demonstrated the practical applicability of DBN in the investigation of intricate biological relationships and regulatory mechanisms. However, traditional DBN learning methods often require centralized data aggregation, which can pose significant privacy risks and scalability challenges, particularly in sensitive domains such as healthcare and genomics. These limitations have spurred interest in privacy-preserving and distributed learning frameworks. Using Federated Learning, DBNs can be trained collaboratively across distributed datasets, enabling secure and scalable model inference while preserving data privacy.
To the best of our knowledge, there is very limited work on federated or privacy-preserving approaches for learning Dynamic Bayesian Networks. However, there are numerous studies based on distributed learning for Bayesian Networks. Gou et al. [16] adopted a two-step procedure that first estimates the BN structures independently using each client’s local dataset and then applies a further conditional independence test. Na and Yang [33] proposed a voting mechanism to select edges that were identified by more than half of the clients. These approaches rely solely on the final graphs that are independently derived from each local dataset, which might result in suboptimal performance due to the limited exchange of information. More importantly, Ng and Zhang [35] proposed a distributed Bayesian learning method based on continuous optimization, using ADMM. Notably, no prior work has tackled distributed learning of DBN.
2 Problem Statement
We consider a scenario with a total of $K$ clients, indexed by $k\in\{1,2,\ldots,K\}$. Each client $k$ possesses its own local dataset consisting of $M$ realizations of a stationary time series. Specifically, for the $k$-th client, we have time series $\{x_{m,t}^{k}\}_{t=0}^{T}$ for $x_{m,t}^{k}\in\mathbb{R}^{d}$, where $m\in\{1,2,\ldots,M\}$ indexes each realization, and $d$ is the number of variables. Due to privacy regulations, these time series cannot be shared directly among clients. For clarity and conciseness, the main text omits the subscript $m$ corresponding to multiple time series, with the complete $m$ notation detailed in the supplementary material §7.1, and consider a generic realization $\{x_{t}^{k}\}_{t=0}^{T}$. In this setting, we assume that the data from all clients follow the same underlying distribution. The extension to non-homogeneous distributions is left for future work (see Sec. 6).
For each client $k$, its dataset (for a single realization) can be written as $\mathcal{D}^{k}=\{x_{t}^{k}\}_{t=p}^{T}$, where $t\in\{p,p+1,\ldots,T\}$ and $p$ denotes the autoregressive order. We model our data using Structural Equation Modeling (SEM) and consider a Vector Autoregressive structure (SVAR) of order $p$:
$(x_{t}^{k})^{\top}=(x_{t}^{k})^{\top}W+(x_{t-1}^{k})^{\top}A_{1}+\dots+(x_{t-p}^{k})^{\top}A_{p}+(u_{t}^{k})^{\top},$ (1)
where:
  1. 1.
    $u_{t}^{k}\sim\mathcal{N}(0,I)$ is a vector of centered noise terms, independent across time and variables;
  2. 2.
    $W$ is a weighted adjacency matrix capturing intra-slice dependencies (edges at the same time step);
  3. 3.
    $A_{i}$, for $i\in\{1,\dots,p\}$, are the coefficient matrices capturing inter-slice dependencies (edges across different time steps).
For each client $k$, the effective sample size is $n_{k}=T+1-p$. Given the collection of datasets $X=\bigcup_{k=1}^{K}\mathcal{D}^{k}$ from all clients, the objective is to recover the true $W$ and $A$ in a privacy-preserving manner. We focus on a setup where clients are motivated to collaborate on learning the DBN structure but are only willing to share minimal information, such as model parameters or estimated graphs, about their local datasets. This setup involves horizontal partitioning of the dataset, which means that, while all local datasets share the same set of variables, the sample data differ between clients. The total sample size, $n$, is the sum of the local sample sizes: $n=\sum_{k=1}^{K}n_{k}$.
3 Federated DBN Inference with ADMM
To implement Federated learning to DBN, our method builds on the DYNOTEARS approach introduced by Pamfil et al. [36], which frames the structural learning of linear Dynamical Bayesian Networks as a continuous constrained optimization problem. By incorporating $\ell_{1}$-norm penalties to encourage sparsity, the optimization problem is formulated as follows:
$\begin{split}\min_{W,A}&\ \ell(W,A)\quad\text{subject to}\quad W\text{ is acyclic},\\ &\ell(W,A)=\frac{1}{2n}\|X_{t}-X_{t}W-X_{(t-p):(t-1)}A\|_{F}^{2}+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}.\end{split}$ (2)
For the acyclicity constraint, following Zheng et al. [55],
$h(W)=\mathrm{tr}\left(e^{W\circ W}\right)-d$
is equal to zero if and only if $W$ is acyclic. Here, $\circ$ represents the Hadamard product (element-wise multiplication) of two matrices. By replacing the acyclicity constraint with the equality constraint $h(W)=0$, the problem can be reformulated as an equality-constrained optimization task.
The ADMM [5] is an optimization algorithm designed to solve convex problems by breaking them into smaller, more manageable subproblems. It is particularly effective for large-scale optimization tasks with complex constraints. By following ADMM framework, we decompose the constrained problem (2) into multiple subproblems and utilize an iterative message-passing approach to converge to the final solution. ADMM proves particularly effective when subproblems have closed-form solutions, which we derive for the first subproblem. To cast problem (2) in an ADMM framework, we reformulate it using local variables $B_{1},\dots,B_{K}\in\mathbb{R}^{d\times d}$ for the intra-slice matrices and $D_{1},\dots,D_{K}\in\mathbb{R}^{d\times d}$ for the inter-slice matrices, along with shared global variables $W\in\mathbb{R}^{d\times d}$ and $A\in\mathbb{R}^{d\times d}$, representing an equivalent formulation:
$\min_{B_{k},D_{k},W,A}\sum_{k=1}^{K}\ell_{k}(B_{k},D_{k})+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}$
subject to $\quad\text{subject to }h(W)=0,$
$\quad B_{k}=W,\quad k=1,2,\ldots,K,$
$\quad D_{k}=A,\quad k=1,2,\ldots,K.$
The local variables $B_{1},\dots,B_{K}$ and $D_{1},\dots,D_{K}$ represent the model parameters specific to each client. Notably, this problem resembles the global variable consensus ADMM framework. The constraints $B_{k}=W$ and $D_{k}=A$ are imposed to enforce consistency, ensuring that the local model parameters across clients remain identical.
Since ADMM combines elements of dual decomposition and the augmented Lagrangian method, making it efficient for handling separable objective functions and facilitating parallel computation, we employ the augmented Lagrangian method to transform the constrained problem into a series of unconstrained subproblems. The augmented Lagrangian is given by:
($\mathcal{L}\left(\{B_{k},D_{k}\}_{k=1}^{K},W,A,\alpha,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\rho_{1},\rho_{2}\right)=$
$\sum_{k=1}^{K}\Bigg{[}\ell_{k}(B_{k},D_{k})+\text{tr}\left(\beta_{k}^{\top}(B_{k}-W)\right)+\frac{\rho_{2}}{2}\|B_{k}-W\|_{F}^{2}$
$\quad+\text{tr}\left(\gamma_{k}^{\top}(D_{k}-A)\right)+\frac{\rho_{2}}{2}\|D_{k}-A\|_{F}^{2}\Bigg{]}$
$+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}+\alpha h(W)+\frac{\rho_{1}}{2}h(W)^{2},$
where $\{\beta_{k},\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$ and $\alpha\in\mathbb{R}$ are estimations of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are the penalty coefficients and $\|*\|_{F}$ is Frobenius norm. Then, we have the iterative update rules of ADMM as follows:
Local Updates for $B_{k}$ and $D_{k}$
$\begin{split}(B_{k}^{(t+1)},D_{k}^{(t+1)})=&\arg\min_{B_{k},D_{k}}\bigg{[}\ell_{k}(B_{k},D_{k})+\text{tr}\left(\beta_{k}^{(t)\top}(B_{k}-W^{(t)})\right)\\ &+\frac{\rho_{2}^{(t)}}{2}\|B_{k}-W^{(t)}\|_{F}^{2}+\text{tr}\left(\gamma_{k}^{(t)\top}(D_{k}-A^{(t)})\right)+\frac{\rho_{2}^{(t)}}{2}\|D_{k}-A^{(t)}\|_{F}^{2}\bigg{]}.\end{split}$ (3)
Global Updates for $W$ and $A$
$\begin{split}(W^{(t+1)},A^{(t+1)})=&\arg\min_{W,A}\Bigg{[}\alpha^{(t)}h(W)+\frac{\rho_{1}^{(t)}}{2}h(W)^{2}+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}\\ &+\sum_{k=1}^{K}\left(\text{tr}\left(\beta_{k}^{(t)\top}(B_{k}^{(t+1)}-W)\right)+\frac{\rho_{2}^{(t)}}{2}\|B_{k}^{(t+1)}-W\|_{F}^{2}\right)\\ &+\sum_{k=1}^{K}\left(\text{tr}\left(\gamma_{k}^{(t)\top}(D_{k}^{(t+1)}-A)\right)+\frac{\rho_{2}^{(t)}}{2}\|D_{k}^{(t+1)}-A\|_{F}^{2}\right)\Bigg{]}.\end{split}$ (4)
Update Dual Variables
$\begin{split}\beta_{k}^{(t+1)}&=\beta_{k}^{(t)}+\rho_{2}^{(t)}(B_{k}^{(t+1)}-W^{(t+1)}),\\ \gamma_{k}^{(t+1)}&=\gamma_{k}^{(t)}+\rho_{2}^{(t)}(D_{k}^{(t+1)}-A^{(t+1)}),\\ \alpha^{(t+1)}&=\alpha^{(t)}+\rho_{1}^{(t)}h(W^{(t+1)}),\\ \rho_{1}^{(t+1)}&=\phi_{1}\rho_{1}^{(t)},\\ \rho_{2}^{(t+1)}&=\phi_{2}\rho_{2}^{(t)},\end{split}$ (5)
where $\phi_{1},\phi_{2}\in\mathbb{R}$ are hyperparameters that control how fast the coefficients $\rho_{1},\rho_{2}$ are increased.
As previously mentioned, ADMM is particularly efficient when the optimization subproblems have closed-form solutions. The subproblem in equation (3) is a well-known proximal minimization problem, extensively explored in the field of numerical optimization [9, 37]. For readability, we define the following matrices:
$S$ $=\frac{1}{n_{k}}X_{t}^{k\top}X_{t}^{k},$
$M$ $=\frac{1}{n_{k}}X_{t}^{k\top}X_{(t-p):(t-1)}^{k},$
$N$ $=\frac{1}{n_{k}}X_{(t-p):(t-1)}^{k\top}X_{(t-p):(t-1)}^{k},$
$P$ $=S+\rho_{2}^{(t)}I,$
$Q$ $=N+\rho_{2}^{(t)}I,$
and vectors:
$b_{1}$ $=S-\beta_{k}^{(t)}+\rho_{2}^{(t)}W^{(t)},$
$b_{2}$ $=M^{\top}-\gamma_{k}^{(t)}+\rho_{2}^{(t)}A^{(t)}.$
By computing the gradient, we can derive the following closed-form solutions:
  • $B_{k}^{(t+1)}$:
    $\boxed{B_{k}^{(t+1)}=\left(P-MQ^{-1}M^{\top}\right)^{-1}\left(b_{1}-MQ^{-1}b_{2}\right).}$
  • $D_{k}^{(t+1)}$:
    $\boxed{D_{k}^{(t+1)}=\left(Q-M^{\top}P^{-1}M\right)^{-1}\left(b_{2}-M^{\top}P^{-1}b_{1}\right).}$
For a comprehensive procedure, we have summarized it in the supplementary material §7.4. The presence of the acyclicity term $h(W)$ prevents us from obtaining a closed-form solution for the problem in equation (4). Instead, the optimization problem can be addressed using first-order methods, such as gradient descent, or second-order methods, like L-BFGS [6]. In this work, we reproduce the approach of Pamfil et al. [36] and employ the L-BFGS method to solve the optimization problem. We have summarized our procedure in Algorithm 1.
Algorithm 1 Distributed DBN Inference with ADMM
0:  Initial parameters $\rho_{1},\rho_{2},\alpha^{(1)},\beta_{1}^{(1)},\ldots,\beta_{K}^{(1)},\gamma_{1}^{(1)},\ldots,\gamma_{K}^{(1)}$; multiplicative factors $\phi_{1},\phi_{2}>1$; initial points $W^{(1)}$ and $A^{(1)}$
1:  for $t=1,2,\ldots$ do
2:     Each client solves problem (3) in parallel
3:     Central server collects $B_{1}^{(t+1)},\ldots,B_{K}^{(t+1)},D_{1}^{(t+1)},\ldots,D_{K}^{(t+1)}$ from all clients
4:     Central server solves problem (4)
5:     Central server sends $W^{(t+1)}$ and $A^{(t+1)}$ to all clients
6:     Central server updates ADMM parameters $\alpha^{(t+1)},\rho_{1}^{(t+1)},\rho_{2}^{(t+1)}$ according to Eq.(5)
7:     Each client updates $\beta_{k}^{(t+1)},\gamma_{k}^{(t+1)}$ according to Eq.(5)
8:  end for
4 Experiments
Figure 1: An example result using 2Dbn for Gaussian noise data with $n=500$ samples, $d=5$ variables, an autoregressive order $p=3$, and $K=10$ clients. We set the thresholds $\tau_{w}=\tau_{a}=0.3$. Our algorithm recovers weights close to the ground truth.
In this section, we first study the performance of 2Dbn on simulated data generated by a linear SVAR structure [15]. We then compare it against three linear baseline methods using three evaluation metrics, demonstrating the effectiveness of our proposed method. Figure 1 provides an illustrative example.
Benchmark Methods.
We compare our ADMM-based approach described in Sec. 3, referred to as 2Dbn, with three other methods. The first is a baseline, denoted as Ave, which computes the average of the weighted adjacency matrices estimated by DYNOTEARS [36] from each client’s local dataset, followed by thresholding to determine the edges. The second baseline, referred to as Best, selects the best graph from among those estimated by each client based on the lowest Structural Hamming Distance (SHD) to the ground truth. While this approach is unrealistic in practical scenarios (since it assumes knowledge of the ground truth), it serves as a useful point of reference. For additional context, we also consider DYNOTEARS applied to the combined dataset from all clients, denoted as Alldata. Note that the final graphs produced by Ave may contain cycles, and we do not apply any post-processing steps to remove them, as doing so could reduce performance. Since there is no official source code for DYNOTEARS, we re-implement it using only the numpy and scipy packages in about a hundred lines of code. This simpler, self-contained implementation eases understanding and reusability compared to available versions on GitHub.
Evaluation Metrics.
We use three metrics to evaluate performance: Structural Hamming Distance, True Positive Rate (TPR), and False Discovery Rate (FDR). SHD measures the dissimilarity between the inferred graph and the ground truth, accounting for missing edges, extra edges, and incorrectly oriented edges [49]. Lower SHD indicates closer alignment with the ground truth. TPR (also known as sensitivity or recall) quantifies the proportion of true edges correctly identified, calculated from true positives and false negatives [13]. Higher TPR means the model identifies more true edges. FDR measures the proportion of false positives among all predicted edges, computed as the ratio of false positives to the sum of false positives and true positives [4]. A lower FDR indicates that most detected edges are correct. Together, these metrics provide a comprehensive understanding of the model’s accuracy in graph structure inference.
Data Generation & Settings.
We generate data following the Structural Equation Model described in Eq. (1). This involves four steps: (1). Constructing weighted graphs $G_{W}$ and $G_{A}$; (2). Creating data matrices $X$ and $Y$ aligned with these graphs; (3). Partitioning the data among $K$ clients as $X_{client}$ and $Y_{client}$; (4). Applying all algorithms to $X_{client}$ and $Y_{client}$ (or $X$ and $Y$) and evaluating their performance. For details of steps (1) and (2), see §7.2. We use Gaussian noise with a standard deviation of 1. The intra-slice DAG is an Erdős-Rényi (ER) graph with a mean degree of 4, and the inter-slice DAG is an ER graph with a mean out-degree of 1. Although this results in very sparse graphs at high $d$, they remain connected under our settings [11]. We set the base of the exponential decay of inter-slice weights to $\eta=1.5$.For the hyperparameters $\lambda_{w}$ and $\lambda_{a}$, we generate heatmaps to systematically identify their optimal values. These values are documented in the Supplementary Materials (§7.3). We set $\phi_{1}=1.6$ and $\phi_{2}=1.1$ with initial $\rho_{1}=\rho_{2}=1$, and the initial Lagrange multipliers are zero. For DYNOTEARS (and thus Ave and Best), we follow the authors’ recommended hyperparameters. To avoid confusion, we summarize the data shape used in 2Dbn:
  • For 2Dbn, $X$ and $Y$ are of size $K\times n^{\prime}\times d$ and $K\times n^{\prime}\times pd$, respectively, where $K$ is the number of clients, $n^{\prime}$ is the number of samples per client, $d$ is the dimensionality, and $p$ is the autoregressive order.
Experiment Settings. We consider two types of experimental settings. First, we fix the number of clients $K$ and increase the number of variables $d$ while maintaining a consistent total sample size $n$. This allows us to assess how well the methods scale with increasing dimensionality. Second, we fix the total sample size $n\in\{256,512\}$ and vary the number of clients $K$ from 2 to 64. This setting evaluates adaptability and robustness as the data becomes more distributed.
4.1 Varying Number of Variables
In this section, we focus on DBNs with $n=5d$ samples evenly distributed across $K=10$ clients for $d=10,20$, and $n=6d$ for $d\in\{5,15\}$. Note that setting $n=5d$ or $6d$ ensures that each client has an integer number of samples. We generate datasets for each of these cases. Typically, each client has very few samples, making this a challenging scenario.
Figure 2: Structure learning results for $W$ in a DBN with Gaussian noise for $d=5,10,15,20$ variables, an autoregressive order $p=1$, and $K=10$ clients.
Figure 2 shows the results for the inferred $W$. The corresponding results for $A$ are provided in §9. Across all tested values of $d$, 2Dbn consistently achieves the lowest SHD compared to Ave and Best, and even outperforms Alldata for $d=5,15$. This advantage may be due to the need for more refined hyperparameter tuning in Alldata. In addition to a lower SHD, 2Dbn achieves a relatively high TPR, close to that of Alldata and higher than that of Best, indicating that 2Dbn accurately identifies most true edges. Moreover, 2Dbn attains a lower FDR than Ave and Best, emphasizing its superior reliability in edge identification under this challenging setup.
4.2 Varying Number of Clients
We now examine scenarios where a fixed total number of samples is distributed among varying numbers of clients. For $d\in\{10,20\}$, we generate $n=512$ samples with $p=1$. These are evenly allocated across $K\in\{2,4,8,16,32,64\}$ clients. For $\lambda_{w}$ and $\lambda_{a}$, we follow the recommendations of [36] for DYNOTEARS, Ave, and Best. For 2Dbn, we select the best regularization parameters from $[0.05,0.5]$ in increments of 0.05 by minimizing SHD.
Figure 3: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients. There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
Figure 4: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients. There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
Figures 3 and 4 show the results for $W$ with $d=10$ and $d=20$. The results for $A$ are in §9. As $K$ increases, the TPRs of Ave and Best drop sharply, resulting in high SHDs. Although 2Dbn’s TPR also decreases as $K$ grows, it remains significantly higher than those of the other baselines. For example, with $d=20$ and $K=64$, 2Dbn achieves a TPR of 0.7, while Ave and Best achieve only 0.5 and 0.3, respectively. This highlights the importance of information exchange in the optimization process, enabling 2Dbn to learn a more accurate DBN structure even in highly distributed scenarios.
Next, we consider the case of $n=256$ samples. The corresponding $W$-based results are shown below, and $A$-based results are in §9.
Figure 5: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients. There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
Figure 6: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients. There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
From Figures 5 and 6, 2Dbn consistently attains the lowest SHD compared with Best and Ave. For $d=20$, the TPRs of Ave and Best continue to decline as $K$ increases, whereas 2Dbn’s TPR remains higher for all $K\geq 8$. At extremely high distributions, such as $K=64$ with $d=20$, each client has only 4 samples, leading to non-convergence for Ave and Best. Thus, we compare metrics at $K=32$ for a fair assessment. For $d=10$, while Ave achieves the highest TPR among all methods for all $K$, its FDR is much larger than that of the other methods. This suggests that the thresholding strategy used by Ave may lead to a high number of false positives. Nonetheless, 2Dbn’s TPR for $d=10$ remains close to that of Alldata as $K$ grows.
In some cases, when $K$ or $d$ is small, the other baselines may outperform 2Dbn. For example, with $d=10,n=512$, Ave and Best achieve better SHDs than 2Dbn when $K\leq 16$. This is expected since fewer clients and lower dimensionality provide each client with enough samples for accurate independent structure learning. However, as complexity grows (e.g., $d=20$), 2Dbn outperforms the baselines by maintaining a lower FDR and effectively leveraging distributed information. These findings highlight the importance of both method selection and hyperparameter tuning, as well as potential directions for future research in improving the efficiency of information exchange mechanisms for DBN learning.
5 Applications
5.1 DREAM4
Building on the DREAM4 Challenge [29, 45, 46], our paper focuses on the time-series track of the InSilico_Size100 subchallenge. In this problem, the DREAM4 gene expression data are used to infer gene regulatory networks. The InSilico_Size100 subchallenge dataset contains 5 independent datasets, each consisting of 10 time-series for 100 genes, measured over 21 time steps. We assume that each dataset corresponds to data collected from different hospitals or research centers, thereby motivating our federated approach. More information and data about DREAM4 are available at https://gnw.sourceforge.net/resources/DREAM4%20in%20silico%20challenge.pdf.
Let $X^{g}_{t,r}$ denote the expression level of gene $g$ at time $t\in\{0,1,2,\ldots,20\}$ in replicate $r\in\{0,1,2,\ldots,R\}$. Note that $R$ depends on the dataset used — for instance, if only one dataset is used, then $R=10$. Consequently, $X_{t,r}\in\mathbb{R}^{100}$ and $X_{t}\in\mathbb{R}^{R\times 100}$. In this experiment, we set $p=1$, aligning with the VAR method proposed in the DREAM4 Challenge by Lu et al. [26]. The federated setting of our 2Dbn approach includes $K=5$ clients, each with $R=2$ replicates. Thus, each client contains a time-series dataset for 100 genes over 42 time steps. We found that small regularization parameters, $\lambda_{w}=\lambda_{a}=0.0025$, work well for all DREAM4 datasets. Lu et al. [26] evaluated various methods for learning these networks, including approaches based on Mutual Information(MI), Granger causality, dynamical systems, Decision Trees, Gaussian Processes (GPs), and Dynamic Bayesian Networks. Notably, we did not compare DYNOTEARS in this study, as its source code is not publicly available, and our implemented version did not achieve optimal results. The comparisons were made using AUPR (Area Under the Precision-Recall Curve) and AUROC (Area Under the Receiver Operating Characteristic Curve) across the 5 datasets. From Table. 1, the GP-based method outperforms all others, achieving the highest mean AUPR (0.208) and AUROC (0.728). In terms of AUPR, our 2Dbn approach outperforms the TSNI (ODE-based) method in its non-federated version. Furthermore, the AUPR of 2Dbn is comparable to those of Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches. Importantly, the mean AUROC of 2Dbn surpasses those of GCCA (VAR-based), ARACNE (MI-based), and TSNI (ODE-based) methods, and is close to those of Ebdnet and VBSSMa (DBN-based).
Overall, 2Dbn is comparable to other non-federated benchmarks. However, the GP-based method still achieves superior results in both AUPR and AUROC. One possible reason for this superiority is that the GP-based approach may more effectively capture nonlinear relationships and complex temporal dynamics, which can be challenging in a federated setting where data distributions and conditions vary across sources. Unsurprisingly, the GP-based method outperforms all approaches because Gaussian Processes naturally handle nonlinear dynamics and are more adaptive to varying, complex relationships. In contrast, the 2Dbn model, although suitable for distributed data, might impose stronger parametric assumptions or face computational constraints that limit its ability to capture subtle nonlinear interactions. Thus, we believe GP-based methods could be extended to a distributed approach for structure learning, making them more applicable to real-world problems. We discuss this further in Section 6.
Method Type AUPR AUROC
2Dbn FED-DBN 0.040 0.600
Ebdnet [41] DBN 0.043 0.640
VBSSMa [39] DBN 0.086 0.620
CSId [38] GP 0.208 0.728
GCCA [39] VAR 0.050 0.584
TSNI [39] ODE 0.026 0.566
ARACNE [30] MI 0.046 0.558
Table 1: Comparison of mean AUPR and mean AUROC scores on the DREAM4 dataset.
5.2 Functional Magnetic Resonance Imaging (FMRI)
In this experiment, we apply the proposed learning methods to estimate connections in the human brain using simulated blood oxygenation level-dependent (BOLD) imaging data [43]. The dataset consists of 28 independent datasets with the number of observed variables $d\in\{5,10,15\}$. Each dataset contains 50 subjects (i.e., 50 ground-truth networks) with 200 time steps. To conduct the experiments, we use simulated time series measurements corresponding to five different human subjects for each $d$ and compute the Average AUROC using the sklearn package.
For the federated setting, we partition the 200 time steps among 5 clients ($K=5$). Detailed information and descriptions of the data are available at https://www.fmrib.ox.ac.uk/datasets/netsim/index.html. In our experiments, we evaluate the proposed method for $d\in\{5,10,15\}$. For $d=5$, we use the 3rd, 6th, 9th, 12th, and 15th subjects from Sim-1.mat. For $d=10$, we use the 2nd, 4th, 6th, 8th, and 10th subjects from Sim-2.mat. Finally, for $d=15$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-3.mat. Further details of this experiment are provided in the supplementary materials (see §7.5). We compare our method to the economy Statistical Recurrent Units (eSRU) proposed by Khanna and Tan [19] for inferring Granger causality, as well as existing methods based on a Multilayer Perceptron (MLP), a Long Short-Term Memory (LSTM) network [48], and a Convolutional Neural Network (CNN)-based model, the Temporal Causal Discovery Framework (TCDF) [34], on multivariate time series data for $d=15$, as examined by Khanna and Tan [19]. As shown in Tab.2, our proposed 2DBNs achieve an AUROC of 0.74, outperforming the LSTM-based approach and approaching the performance of the CNN-based TCDF method. Even though 2DBNs do not surpass the MLP-based and eSRU-based methods, it is notable that the Federated version of our approach outperforms or closely matches several established non-Federated benchmarks. This outcome is reasonable, as the MLP and eSRU methods rely on deep architectures adept at modeling complex structural dependencies. Notably, our 2DBN method provides a new perspective on this problem by ensuring data security through its Federated approach. This capability is particularly important in scenarios involving sensitive or distributed datasets, as it allows for effective analysis without compromising privacy or data integrity. We have further elaborated on this aspect in Sec.6.
Method Averaged AUROC
MLP [48] 0.81$\pm$0.04
LSTM [48] 0.70$\pm$0.03
TCDF [34] 0.75$\pm$0.04
eSRU [19] 0.84$\pm$0.03
2DBNs 0.74$\pm$0.04
Table 2: Mean AUROC comparison of different methods for $d=15$.
6 Discussion
We proposed a federated framework for learning DBNs on horizontally partitioned data. Specifically, we designed a distributed DBN learning method using ADMM, where only model parameters are exchanged during optimization. Our experimental results demonstrate that this approach outperforms other state-of-the-art methods, particularly in scenarios with numerous clients, each possessing a small sample size — a common situation in federated learning that motivates client collaboration. Below, we address some limitations of our approach and suggest potential directions for future research.
Assumptions.
We assumed that the structure of the DBN is fixed over time and is identical for all time series in the dataset (i.e., it is the same for all $m\in M$). Relaxing these assumptions could be useful in various ways, such as allowing the structure to change smoothly over time [44]. Another direction for future work is to investigate the behavior of the algorithm on non-stationary or cointegrated time series [28] or in scenarios with confounding variables [17].
Federated Learning.
The proposed ADMM-based approach to federated learning relies on a stateful setup, requiring each participating client to be involved in every round of communication and optimization. This “always-on” requirement can be burdensome in real-world scenarios. For instance, in large-scale deployments, clients such as mobile devices or IoT sensors may experience intermittent connectivity, limited power, or varying levels of availability. Ensuring that all such devices participate consistently and synchronously in every round is often impractical and can result in significant performance bottlenecks. An important future direction is to explore asynchronous techniques, which would enable stateless clients and facilitate cross-device learning. Furthermore, data may be vertically partitioned across clients, meaning that each client owns different variables but collectively aims to perform Bayesian Network Structure Learning (BNSL). Developing a federated approach tailored to this vertical setting represents another promising research direction. Additionally, the ADMM procedure involves sharing model parameters with a central server, raising concerns about potential privacy risks. Research has shown that these parameters can leak sensitive information in certain scenarios, such as with image data [40]. To address this, exploring differential privacy techniques [10] to enhance the protection of shared model parameters is a critical avenue for future work. Moreover, our current work focuses on non-heterogeneous data. However, heterogeneous settings are more applicable to many real-world problems, as clients often differ in computational capabilities, communication bandwidth, and local data distributions. These variations pose significant challenges for model convergence and performance. Several methods have been proposed to address these issues. For example, FedProx [24] incorporates a proximal term to stabilize and unify updates from clients with varying local training conditions, while FedNova [50] normalizes aggregated updates to mitigate the impact of heterogeneous local computational workloads.
Nonlinear Dependencies.
Finally, we emphasize that the linear assumption of our methodology was made purely for simplicity, allowing us to focus on the most salient dynamic and temporal aspects of the problem. It is possible to model more complex nonlinear dependencies using Gaussian Processes [18, 42, 14, 53, 25] or neural networks. Additionally, the least squares loss function can be replaced with logistic loss (or more generally, any exponential family log-likelihood) to model binary data. Further, it would be valuable to consider combinations of continuous and discrete data [3], which are essential for many real-world applications.

### v2

@@
-Temporal Causal Discovery in Dynamic Bayesian Networks Using Federated Learning
+Federated Learning of Dynamic Bayesian Network via Continuous Optimization from Time Series Data
 Abstract
-Traditionally, learning the structure of a Dynamic Bayesian Network has been centralized, with all data pooled in one location.
-However, in real-world scenarios, data are often dispersed among multiple parties (e.g., companies, devices) that aim to collaboratively learn a Dynamic Bayesian Network while preserving their data privacy and security.
-In this study, we introduce a federated learning approach for estimating the structure of a Dynamic Bayesian Network from data distributed horizontally across different parties.
-We propose a distributed structure learning method that leverages continuous optimization so that only model parameters are exchanged during optimization.
-Experimental results on synthetic and real datasets reveal that our method outperforms other state-of-the-art techniques, particularly when there are many clients with limited individual sample sizes.
+Traditionally, learning the structure of a Dynamic Bayesian Network has been centralized, requiring all data to be pooled in one location.
+However, in real-world scenarios, data are often distributed across multiple entities (e.g., companies, devices) that seek to collaboratively learn a Dynamic Bayesian Network while preserving data privacy and security.
+More importantly, due to the presence of diverse clients, the data may follow different distributions, resulting in data heterogeneity.
+This heterogeneity poses additional challenges for centralized approaches.
+In this study, we first introduce a federated learning approach for estimating the structure of a Dynamic Bayesian Network from homogeneous time series data that are horizontally distributed across different parties.
+We then extend this approach to heterogeneous time series data by incorporating a proximal operator as a regularization term in a personalized federated learning framework.
+To this end, we propose FDBNL and PFDBNL, which leverage continuous optimization, ensuring that only model parameters are exchanged during the optimization process.
+Experimental results on synthetic and real-world datasets demonstrate that our method outperforms state-of-the-art techniques, particularly in scenarios with many clients and limited individual sample sizes.
 1 Introduction
-The learning of Dynamic Bayesian Network (DBN) structures is a crucial technique in causal learning and representation learning.
-More importantly, it has been applied in various real-world applications, such as modeling gene expression [21], measuring network security [8], studying manufacturing process parameters [47], and recognizing interaction activities [54].
-Traditionally, the learning of DBNs has been conducted at a centralized location where all data is aggregated.
-However, with the rapid advancements in technology and the Internet of Things (IoT) [27] devices, data collection has become significantly more accessible.
-Consequently, in real-world scenarios, data is typically owned and dispersed among multiple entities, including mobile devices, individuals, companies, and healthcare institutions.
-In many cases, these entities, referred to as clients, may lack an adequate number of samples to independently construct a meaningful DBN.
-They might seek to collaboratively gain comprehensive insights into the DBN structure without disclosing their individual data due to privacy and security concerns.
-For instance, consider several hospitals aiming to work together to develop a DBN that uncovers the conditional independence relationships among their variables of interest, such as various medical conditions.
-Sharing patients’ records is clearly prohibited by privacy regulations [2].
-Consequently, the sensitive nature of the data renders the centralized aggregation of data from different clients for DBN learning infeasible.
+Learning the structure of a Dynamic Bayesian Network (DBN) is a fundamental technique in causal learning and representation learning.
+DBNs are particularly valuable for uncovering temporal dependencies and conditional independence relationships in time-series data, making them a cornerstone of causal inference in dynamic systems.
+More importantly, DBNs have been widely applied in real-world domains, such as modeling gene expression [26], assessing network security [10], analyzing manufacturing process parameters [55], and recognizing interaction activities [62].
+Traditionally, DBN structure learning has been conducted in a centralized setting, where all data is aggregated in one location.
+However, with the rapid advancements in technology and the proliferation of Internet of Things (IoT) [31] devices, data collection has become significantly more decentralized.
+In real-world scenarios, data is often owned and distributed across multiple entities, such as mobile devices, individuals, companies, and healthcare institutions.
+Many of these entities, referred to as clients, may lack a sufficient number of samples to independently construct a meaningful DBN.
+To gain comprehensive insights, they may seek to collaboratively learn the DBN structure while preserving data privacy and security.
+For instance, consider several hospitals aiming to collaboratively develop a DBN to uncover conditional independence relationships among medical conditions.
+Sharing patient records is strictly prohibited by privacy regulations [2], making centralized data aggregation infeasible.
+Furthermore, data heterogeneity introduces additional complexities: clients may exhibit homogeneous data (similar statistical distributions) or heterogeneous data, where distributions differ significantly or feature spaces (e.g., variables or attributes) vary entirely.
+These disparities complicate collaborative learning, as models must reconcile mismatched patterns or structural differences without direct data sharing.
 A robust privacy-preserving learning strategy is therefore essential.
-In recent years, federated learning (FL) [31] has gained significant attention, enabling numerous clients to collaboratively train a machine learning model through a coordinated and privacy-preserving approach.
-In this framework, each client is not required to share its raw data; instead, they only provide the minimal necessary information, such as model parameters and gradient updates, pertinent to the specific learning task.
-This approach has been applied across various research domains, including the Internet of Things [20] and recommender systems [7].
-Readers may refer to these review papers for more comprehensive insights [23, 51].
-Traditionally, score-based learning of DBNs relies on discrete optimization, making it incompatible with common continuous optimization approaches used in federated learning methods.
-Fortunately, recent work by Zheng et al.
+In recent years, federated learning (FL) [35] has gained significant attention as a framework that enables multiple clients to collaboratively train machine learning models in a privacy-preserving manner.
+In FL, clients do not share raw data but instead exchange minimal necessary information, such as model parameters and gradient updates, relevant to the learning task.
+This approach has been successfully applied across various domains, including the Internet of Things [25], condition monitoring [4, 63], and recommender systems [8].
+Readers seeking further insights into FL can refer to comprehensive review papers [28, 59].
+Moreover, Personalized Federated Learning (PFL) extends FL to address the challenges posed by heterogeneous data distributions across clients.
+Unlike traditional FL, which trains a single global model, PFL allows for the development of personalized models tailored to each client’s unique data distribution while still benefiting from collaborative learning.
+To achieve this balance between personalization and collaboration, various PFL approaches have been proposed, including proximal regularization [48] to balance local and global models, clustered federated learning [16, 49, 36] to group clients with similar data distributions, and meta-learning-based methods [22, 9] to facilitate fast local adaptation.
+These techniques enhance performance in real-world federated environments while preserving data privacy and minimizing communication costs.
+Traditionally, score-based DBN learning relies on discrete optimization, making it incompatible with common continuous optimization techniques.
+However, continuous optimization methods, particularly those used in federated learning, solve problems efficiently using convex optimization techniques.
+Recent works by Zheng et al.
 [64] and Pamfil et al.
-[36] have made continuous optimization strategies possible for DBN learning by providing algebraic characterizations of acyclicity.
-Despite this advancement, directly applying popular federated optimization techniques (FedAvg [31], Ditto [22]) to these continuous formulations remains challenging due to the acyclicity constraint inherent in DBN structure learning.
+[41] have addressed this limitation by introducing continuous optimization strategies for DBN learning, leveraging algebraic characterizations of acyclicity.
+Despite this advancement, directly applying popular federated optimization techniques - such as FedAvg [35] and Ditto [27] - to these continuous formulations remains challenging due to the inherent acyclicity constraint in DBN structure learning.
 Contributions.
-In this work, we present a federated learning approach for estimating the structure of Dynamic Bayesian Network from observational data (with temporal information) that is horizontally partitioned across different agents.
-Our contributions are threefold.
-  • We propose a Federated DBN Learning method 2Dbn based on continuous optimization using the Alternating Direction Method of Multipliers (ADMM).
-2Dbn can simultaneously learn contemporaneous and time-lagged causal structure without any implicit assumptions on the underlying graph topology.
-  • Using extensive simulations and two real-world, high-dimensional datasets, we demonstrate that 2Dbn is scalable and effectively identifies the true underlying DBN graph.
-  • To our knowledge, this is the first study to explore learning DBN within the FL framework.
-Our full code is available on GitHub: https://github.com/PeChen123/2DBN_learning.
-Literature Review.
-Dynamic Bayesian Network are probabilistic models that extend static Bayesian Networks to represent multi-stage processes by modeling sequences of variables over different stages.
-Learning DBN involves two key tasks: structure learning (identifying the network topology) and parameter learning (estimating the conditional probabilities).
+In this work, we introduce a federated learning approach and a personalized federated learning approach for estimating the structure of a Dynamic Bayesian Network from homogeneous and heterogeneous observational time-series data that is horizontally partitioned across different clients.
+Our contributions can be summarized as follows:
+  • Novel Federated Approach for DBN Learning:
+    • Federated DBN Learning (FDBNL) for homogeneous time series, formulated as a continuous optimization problem and solved using the Alternating Direction Method of Multipliers (ADMM).
+    • Personalized Federated DBN Learning (PFDBNL) for heterogeneous time series, incorporating a proximal operator to account for client-specific variations.
+  • Both FDBNL and PFDBNL enable causal structure learning directly from data without imposing restrictive assumptions on the underlying graph topology.
+  • Through extensive simulations and experiments on two high-dimensional, real-world datasets, we demonstrate that our methods are scalable and effectively recover the true underlying DBN structures, even in challenging heterogeneous settings.
+  • To the best of our knowledge, this is the first study to explore DBN structure learning within the federated learning framework.
+Additionally, we pioneer the study of temporal dynamic structure learning from heterogeneous data in a federated context.
+Our full code is available on GitHub: https://github.com/PeChen123/FDBN_Learning.
+Literature Review
+Dynamic Bayesian Networks are probabilistic models that extend static Bayesian Networks to represent multi-stage processes by modeling sequences of variables across different time steps.
+Learning a DBN involves two key tasks: structure learning (identifying the network topology) and parameter learning (estimating the conditional probabilities).
 Friedman et al.
@@
 Building on this work, Murphy [37] adapted static Bayesian Network learning techniques for temporal domains, employing methods such as the Expectation-Maximization algorithm for parameter estimation.
-Advancements in structure learning from time series data were further propelled by Pamfil et al.
+Advancements in structure learning from time-series data were further propelled by Pamfil et al.
 [41], who developed an extension of Bayesian Network inference using continuous optimization methods.
 Their approach enabled efficient and scalable learning of DBN structures by formulating the problem as a continuous optimization task with acyclicity constraints tailored for temporal graphs.
-To address non-linear dependencies in time series data, Tank et al.
+To address non-linear dependencies in time-series data, Tank et al.
 [56] introduced neural Granger causality methods, which leverage neural networks within the DBN framework to capture complex non-linear interactions.
@@
 [60] enhanced Bayesian Network inference techniques to generate causal networks from observational biological data.
-Their work demonstrated the practical applicability of DBN in the investigation of intricate biological relationships and regulatory mechanisms.
-However, traditional DBN learning methods often require centralized data aggregation, which can pose significant privacy risks and scalability challenges, particularly in sensitive domains such as healthcare and genomics.
+Their work demonstrated the practical applicability of DBNs in investigating intricate biological relationships and regulatory mechanisms.
+However, traditional DBN learning methods often require centralized data aggregation, which poses significant privacy risks and scalability challenges, particularly in sensitive domains such as healthcare and genomics.
 These limitations have spurred interest in privacy-preserving and distributed learning frameworks.
-Using Federated Learning, DBNs can be trained collaboratively across distributed datasets, enabling secure and scalable model inference while preserving data privacy.
+Federated Learning enables DBNs to be trained collaboratively across distributed datasets, facilitating secure and scalable model inference while preserving data privacy.
 To the best of our knowledge, there is very limited work on federated or privacy-preserving approaches for learning Dynamic Bayesian Networks.
-However, there are numerous studies based on distributed learning for Bayesian Networks.
+However, there are numerous studies on distributed learning for Bayesian Networks.
 Gou et al.
@@
 Na and Yang [38] proposed a voting mechanism to select edges that were identified by more than half of the clients.
-These approaches rely solely on the final graphs that are independently derived from each local dataset, which might result in suboptimal performance due to the limited exchange of information.
-More importantly, Ng and Zhang [35] proposed a distributed Bayesian learning method based on continuous optimization, using ADMM.
-Notably, no prior work has tackled distributed learning of DBN.
+These approaches rely solely on the final graphs independently derived from each local dataset, which may lead to suboptimal performance due to the limited exchange of information.
+More importantly, Ng and Zhang [40] proposed a distributed Bayesian Network learning method based on continuous optimization using the Alternating Direction Method of Multipliers (ADMM).
+Notably, no prior work has addressed distributed learning of DBNs.
 2 Problem Statement
-We consider a scenario with a total of $K$ clients, indexed by $k\in\{1,2,\ldots,K\}$.
-Each client $k$ possesses its own local dataset consisting of $M$ realizations of a stationary time series.
-Specifically, for the $k$-th client, we have time series $\{x_{m,t}^{k}\}_{t=0}^{T}$ for $x_{m,t}^{k}\in\mathbb{R}^{d}$, where $m\in\{1,2,\ldots,M\}$ indexes each realization, and $d$ is the number of variables.
-Due to privacy regulations, these time series cannot be shared directly among clients.
-For clarity and conciseness, the main text omits the subscript $m$ corresponding to multiple time series, with the complete $m$ notation detailed in the supplementary material §7.1, and consider a generic realization $\{x_{t}^{k}\}_{t=0}^{T}$.
-In this setting, we assume that the data from all clients follow the same underlying distribution.
-The extension to non-homogeneous distributions is left for future work (see Sec.
-6).
-For each client $k$, its dataset (for a single realization) can be written as $\mathcal{D}^{k}=\{x_{t}^{k}\}_{t=p}^{T}$, where $t\in\{p,p+1,\ldots,T\}$ and $p$ denotes the autoregressive order.
-We model our data using Structural Equation Modeling (SEM) and consider a Vector Autoregressive structure (SVAR) of order $p$:
-$(x_{t}^{k})^{\top}=(x_{t}^{k})^{\top}W+(x_{t-1}^{k})^{\top}A_{1}+\dots+(x_{t-p}^{k})^{\top}A_{p}+(u_{t}^{k})^{\top},$ (1)
+We consider a scenario with $K$ clients, indexed by $k\in\{1,2,\ldots,K\}$.
+Each client $k$ possesses a local dataset containing $M$ realizations of a stationary time series.
+Specifically, for the $k$-th client, the time series dataset is denoted as $\mathcal{D}^{k}=\{x_{m,t}^{k}\}_{t=0}^{T}$, where $x_{m,t}^{k}\in\mathbb{R}^{d}$, $m\in\{1,2,\ldots,M\}$ indexes realizations, and $d$ is the number of variables.
+Due to privacy constraints, these time series cannot be directly shared between clients.
+For simplicity, we omit the realization subscript $m$ in the main text (see supplementary material §7.1 for full notation) and consider a generic realization $\{x_{t}^{k}\}_{t=0}^{T}$.
+For each client $k$, we model the data $\{x_{t}^{k}\}$ (for a single realization) using a Structural Equation Model (SEM) with a Structural Vector Autoregressive (SVAR) structure of order $p$:
+$(x_{t}^{k})^{\top}=(x_{t}^{k})^{\top}W_{k}+(x_{t-1}^{k})^{\top}A_{k_{1}}+\dots+(x_{t-p}^{k})^{\top}A_{k_{p}}+(u_{t}^{k})^{\top},$
 where:
   1. 1.
-    $u_{t}^{k}\sim\mathcal{N}(0,I)$ is a vector of centered noise terms, independent across time and variables;
+    $t\in\{p,p+1,\ldots,T\}$, and $p$ is the autoregressive order.
   2. 2.
-    $W$ is a weighted adjacency matrix capturing intra-slice dependencies (edges at the same time step);
+    $u_{t}^{k}\sim\mathcal{N}(0,I)$ is a noise vector, independent across time and variables.
   3. 3.
-    $A_{i}$, for $i\in\{1,\dots,p\}$, are the coefficient matrices capturing inter-slice dependencies (edges across different time steps).
-For each client $k$, the effective sample size is $n_{k}=T+1-p$.
-Given the collection of datasets $X=\bigcup_{k=1}^{K}\mathcal{D}^{k}$ from all clients, the objective is to recover the true $W$ and $A$ in a privacy-preserving manner.
-We focus on a setup where clients are motivated to collaborate on learning the DBN structure but are only willing to share minimal information, such as model parameters or estimated graphs, about their local datasets.
-This setup involves horizontal partitioning of the dataset, which means that, while all local datasets share the same set of variables, the sample data differ between clients.
-The total sample size, $n$, is the sum of the local sample sizes: $n=\sum_{k=1}^{K}n_{k}$.
-3 Federated DBN Inference with ADMM
-To implement Federated learning to DBN, our method builds on the DYNOTEARS approach introduced by Pamfil et al.
-[36], which frames the structural learning of linear Dynamical Bayesian Networks as a continuous constrained optimization problem.
+    $W_{k}$ is a weighted adjacency matrix (structured as a directed acyclic graph (DAG)) capturing intra-slice dependencies for the $k$-th client.
+  4. 4.
+    $A_{k_{i}}$ ($i\in\{1,\dots,p\}$) are coefficient matrices capturing inter-slice dependencies for the $k$-th client.
+The matrices $W_{k}$ and $A_{k_{i}}$ may be homogeneous (shared across all clients) or heterogeneous (client-specific).
+We assume the order $p$ is same for all clients.
+In the homogeneous case, we write $W_{k}=W$ and $A_{k_{i}}=A_{i}$ since $W_{k}$ and $A_{k_{i}}$ are the same for all clients.
+In the heterogeneous case, these matrices vary by client.
+The sample size for client $k$ is $n_{k}=T+1-p$.
+Given the federated dataset $X=\bigcup_{k=1}^{K}\mathcal{D}^{k}$, the goal is to recover the true parameters $W,A_{i}$ (homogeneous) or $W_{k},A_{k_{i}}$ (heterogeneous) in a privacy-preserving manner.
+Clients collaborate to learn the DBN structure, combining intra-slice ($W_{k}$) and inter-slice ($A_{k_{i}}$) dependencies, while sharing only minimal information (e.g., model parameters or graphs).
+The total sample size is $n=\sum_{k=1}^{K}n_{k}$.
+We also consider partial client participation, where only a subset of clients $j\leq K$ contributes in each communication round.
+This is critical for scalability, as inactive clients can rejoin later without disrupting the process.
+Notably, partial participation is primarily relevant in the heterogeneous case, where client-specific parameters require updates from participating clients.
+In the homogeneous case, shared parameters remain unaffected by non-participating clients.
+3 Proposed Methodologies
+In this section, we introduce the framework for Federated Dynamic Bayesian Network Learning and Personalized Federated Dynamic Bayesian Network Learning. The general idea is summarized in Figure 1.
+Figure 1: Overview of Personalized Federated DBN learning with $d=3$ nodes, autoregression order $p=2$, and $K=3$ clients.
+The only difference between it and Federated DBN learning is that Federated DBN learning requires each $W_{k}$ and $A_{k_{i}}$ to be identical across clients.
+3.1 Federated Dynamic Bayesian Network Learning
+In this section, we introduce the Federated DBN learning (FDBNL) for the homogeneous (i.i.d) time series.
+To implement federated learning to DBN, our method builds on the method introduced by Pamfil et al.
+[41], which frames the structural learning of linear DBN as a continuous constrained optimization problem.
 By incorporating $\ell_{1}$-norm penalties to encourage sparsity, the optimization problem is formulated as follows:
-$\begin{split}\min_{W,A}&\ \ell(W,A)\quad\text{subject to}\quad W\text{ is acyclic},\\ &\ell(W,A)=\frac{1}{2n}\|X_{t}-X_{t}W-X_{(t-p):(t-1)}A\|_{F}^{2}+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}.\end{split}$ (2)
-For the acyclicity constraint, following Zheng et al. [55],
+$\begin{split}\min_{W,A}&\ \ell(W,A)+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}\quad\text{subject to}\quad W\text{ is acyclic},\\ &\ell(W,A)=\frac{1}{2n}\|X_{t}-X_{t}W-X_{(t-p):(t-1)}A\|_{F}^{2},\end{split}$ (1)
+where $X_{(t-p):(t-1)}=\left[X_{t-1}\ \cdots\ X_{t-p}\right]$, with each $X_{t-i}$ (for $i\in\{1,\ldots,p\}$) being a time-lagged version of $X_{t}$ and $n\times d$ matrix where rows correspond to realizations.
+The matrix $A$ is defined as:
+$A=\begin{bmatrix}A_{1}^{\top}\\ \vdots\\ A_{p}^{\top}\end{bmatrix},$
+where each $A_{i}\in\mathbb{R}^{d\times d}$ captures lag-$i$ dependencies.
+Thus, we observe that $X_{(t-p):(t-1)}$ is an $n\times(p\cdot d)$ matrix, while $A$ is a $(p\cdot d)\times d$ matrix.
+For the acyclicity constraint, following Zheng et al.
+[64],
 $h(W)=\mathrm{tr}\left(e^{W\circ W}\right)-d$
@@
 By replacing the acyclicity constraint with the equality constraint $h(W)=0$, the problem can be reformulated as an equality-constrained optimization task.
-The ADMM [5] is an optimization algorithm designed to solve convex problems by breaking them into smaller, more manageable subproblems.
+This setting does not work for federated learning since all data is encapsulated.
+Thus, for the federated setting, we leverage the Alternating Direction Method of Multipliers [6], an optimization algorithm designed to solve convex problems by decomposing them into smaller, more manageable subproblems.
 It is particularly effective for large-scale optimization tasks with complex constraints.
@@
 ADMM proves particularly effective when subproblems have closed-form solutions, which we derive for the first subproblem.
-To cast problem (2) in an ADMM framework, we reformulate it using local variables $B_{1},\dots,B_{K}\in\mathbb{R}^{d\times d}$ for the intra-slice matrices and $D_{1},\dots,D_{K}\in\mathbb{R}^{d\times d}$ for the inter-slice matrices, along with shared global variables $W\in\mathbb{R}^{d\times d}$ and $A\in\mathbb{R}^{d\times d}$, representing an equivalent formulation:
+To cast problem (1) in an ADMM framework, we reformulate it using local variables $B_{1},\dots,B_{K}\in\mathbb{R}^{d\times d}$ for the intra-slice matrices and $D_{1},\dots,D_{K}\in\mathbb{R}^{pd\times d}$ for the inter-slice matrices, along with shared global variables $W\in\mathbb{R}^{d\times d}$ and $A\in\mathbb{R}^{pd\times d}$, representing an equivalent formulation:
 $\min_{B_{k},D_{k},W,A}\sum_{k=1}^{K}\ell_{k}(B_{k},D_{k})+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}$
@@
 The local variables $B_{1},\dots,B_{K}$ and $D_{1},\dots,D_{K}$ represent the model parameters specific to each client.
+The data for the $k$-th client consists $X_{t}^{k}$ and $X^{k}_{(t-p):(t-1)}$.
 Notably, this problem resembles the global variable consensus ADMM framework.
@@
 The augmented Lagrangian is given by:
-($\mathcal{L}\left(\{B_{k},D_{k}\}_{k=1}^{K},W,A,\alpha,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\rho_{1},\rho_{2}\right)=$
-$\sum_{k=1}^{K}\Bigg{[}\ell_{k}(B_{k},D_{k})+\text{tr}\left(\beta_{k}^{\top}(B_{k}-W)\right)+\frac{\rho_{2}}{2}\|B_{k}-W\|_{F}^{2}$
+($\mathcal{L}\left(\{B_{k},D_{k}\}_{k=1}^{K},W,A,\alpha,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\rho_{1},\rho_{2}\right)$ $=\sum_{k=1}^{K}\Bigg{[}\ell_{k}(B_{k},D_{k})+\text{tr}\left(\beta_{k}^{\top}(B_{k}-W)\right)+\frac{\rho_{2}}{2}\|B_{k}-W\|_{F}^{2}$
 $\quad+\text{tr}\left(\gamma_{k}^{\top}(D_{k}-A)\right)+\frac{\rho_{2}}{2}\|D_{k}-A\|_{F}^{2}\Bigg{]}$
 $+\lambda_{W}\|W\|_{1}+\lambda_{A}\|A\|_{1}+\alpha h(W)+\frac{\rho_{1}}{2}h(W)^{2},$
-where $\{\beta_{k},\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$ and $\alpha\in\mathbb{R}$ are estimations of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are the penalty coefficients and $\|*\|_{F}$ is Frobenius norm.
+where $\{\beta_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$, $\{\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{pd\times d}$ and $\alpha\in\mathbb{R}$ are estimations of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are the penalty coefficients and $\|\cdot\|_{F}$ is Frobenius norm.
 Then, we have the iterative update rules of ADMM as follows:
@@
 Instead, the optimization problem can be addressed using first-order methods, such as gradient descent, or second-order methods, like L-BFGS [7].
-In this work, we reproduce the approach of Pamfil et al.
-[36] and employ the L-BFGS method to solve the optimization problem.
+In this work, we write $W=W_{+}-W_{-}$ (and analogously for $A$) to handle the $\ell_{1}$ penalty term and employ the L-BFGS method to solve the optimization problem.
 We have summarized our procedure in Algorithm 1.
-Algorithm 1 Distributed DBN Inference with ADMM
+Algorithm 1 Federated Dynamic Bayesian Network Learning
 0:  Initial parameters $\rho_{1},\rho_{2},\alpha^{(1)},\beta_{1}^{(1)},\ldots,\beta_{K}^{(1)},\gamma_{1}^{(1)},\ldots,\gamma_{K}^{(1)}$; multiplicative factors $\phi_{1},\phi_{2}>1$; initial points $W^{(1)}$ and $A^{(1)}$
@@
 8:  end for
+3.2 Personalized Federated Dynamic Bayesian Network Learning
+In this section, we introduce Personalized Federated DBN Learning (PFDBNL) for heterogeneous time series data.
+This method is based on the proximal operator, which is used to reformulate Problem.
+(1), as introduced by [43].
+The formal definition of the proximal operator is:
+Definition 1 (Proximal Operator)
+Let $\mathcal{X}$ denote a vector space with norm .$\|.\|_{\mathcal{X}}$, The proximal operator is defined as
+$prox_{\mu f}(X)=\arg\min_{Z}{\mu||Z-X||_{\mathcal{X}}^{2}+f(Z)}$
+To formulate the objective function for PFDBNL, we let
+$f_{k}(W_{k},A_{k})=\ell_{k}(W_{k},A_{k})+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}$ (5)
+for the $k$-th client.
+Note that we sometimes write $A_{k}=\left[A^{\top}_{k_{1}}\ \cdots\ A^{\top}_{k_{p}}\right]^{\top}$ for PFDBNL-related representations.
+Then, we adopt the method described in [48], where $f_{k}$ is replaced with its proximal operator.
+The detailed structure of the resulting optimization problem is outlined as follows:
+$\begin{split}F_{k}(W,A)\coloneqq prox_{\mu f}(W,A)&=\min_{W_{k},A_{k}}\mu||W_{k}-W||^{2}_{F}+\mu||A_{k}-A||^{2}_{F}+f_{k}(W_{k},A_{k})\\ &=\min_{W_{k},A_{k}}\ell(W_{k},A_{k})+\mu||W_{k}-W||^{2}_{F}+\mu||A_{k}-A||^{2}_{F}+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}.\end{split}$ (6)
+The hyperparameter $\mu$ regulates the balance between the global model $(W,A)$ and the personalized model $(W_{k},A_{k})$.
+A larger $\mu$ benefits clients with less reliable data by leveraging aggregated information from other clients, while a smaller $\mu$ prioritizes personalization for clients that possess a significant amount of high-quality data.
+Therefore, we define the following optimization problem:
+$\begin{split}\min_{W,A}&\ \sum_{k=1}^{K}F_{k}(W,A)\quad\text{subject to}\quad W\text{ is acyclic},\\ &F_{k}(W,A)=\min_{W_{k},A_{k}}\ell(W_{k},A_{k})+\mu||W_{k}-W||^{2}_{F}+\mu||A_{k}-A||^{2}_{F}+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}.\end{split}$ (7)
+Problem (7) forms a bi-level optimization problem, which is traditionally solved by iteratively applying first-order gradient methods.
+The lower-level problem is solved first to approximate $W_{i}$, and this solution is then used in the upper-level problem to optimize $W$, repeating the process until convergence.
+However, while this approach is straightforward, it often suffers from low solution accuracy and challenges in fine-tuning parameters such as the learning rate.
+To overcome these limitations, we propose a relaxed formulation of Problem (7) to improve both efficiency and accuracy.
+$\begin{split}&\min_{\{W_{k},A_{k}\}_{k=1}^{K},W,A,}\sum_{k=1}^{K}\ell_{k}(W_{k},A_{k})+\mu(||W_{k}-W||^{2}_{F}+||A_{k}-A||^{2}_{F})+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}\\ &\quad\text{subject to }h(W_{k})=0.\\ \end{split}$ (8)
+Our goal is to learn an optimal personalized model $(W_{k},A_{k})$ for each client $k$ and an optimal global model $(W,A)$ that jointly minimize Problem (8), ensuring a balance between personalization and global knowledge aggregation.
+Since ADMM is a primal-dual method known for its improved iteration stability and faster convergence compared to gradient-based methods, we propose using ADMM to solve Problem (8).
+To facilitate this, we introduce auxiliary variables $\{\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K}$, allowing us to reformulate Problem (8) into a separable structure by partitioning the variables into multiple blocks.
+Specifically, this transformation enables a more efficient optimization framework, as detailed below:
+$\begin{split}&\min_{\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K},W,A,}\sum_{k=1}^{K}\ell_{k}(W_{k},A_{k})+\mu(||W_{k}-\tilde{W}_{k}||^{2}_{F}+||A_{k}-\tilde{A}_{k}||^{2}_{F})+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}\\ &\quad\text{subject to }h(W_{k})=0,\\ &\quad\tilde{W}_{k}=W\\ &\quad\tilde{A}_{k}=A,\quad k=1,2,\ldots,K.\end{split}$ (9)
+Here, $\{\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K}$ represents the local model of client $k$.
+Problem (9) remains equivalent to Problem (8) in the sense that they share the same optimal solutions.
+Since Problem (9) involves linear constraints and multiple block variables, we can solve it by employing the exact penalty method within the ADMM framework, which is well-suited for handling such optimization structures.
+Furthermore, we assume $h(W_{k})=0$ for all clients, as we need to enforce the DAG constraint across all clients.
+Similarly, we apply the augmented Lagrangian method to reformulate the problem into a series of unconstrained subproblems.
+To implement ADMM for this problem, we define the corresponding augmented Lagrangian function as follows:
+($\mathcal{L}\Bigl{(}\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K},\,W,\,A,\,\{\beta_{k},\gamma_{k}\}_{k=1}^{K};\,\alpha,\,\rho_{1},\,\rho_{2},\,\mu\Bigr{)}=$
+$\quad\sum_{k=1}^{K}\Bigl{[}\ell_{k}\bigl{(}W_{k},A_{k}\bigr{)}\;+\;\mu\,\|W_{k}-\tilde{W}_{k}\|_{F}^{2}\;+\;\mu\,\|A_{k}-\tilde{A}_{k}\|_{F}^{2}\;+\;\lambda_{W}\,\|W_{k}\|_{1}\;+\;\lambda_{A}\,\|A_{k}\|_{1}$
+$\qquad\quad+\;\mathrm{tr}\Bigl{(}\beta_{k}^{\top}\!(\tilde{W}_{k}-W)\Bigr{)}\;+\;\frac{\rho_{2}}{2}\,\|\tilde{W}_{k}-W\|_{F}^{2}\;+\;\mathrm{tr}\Bigl{(}\gamma_{k}^{\top}\!(\tilde{A}_{k}-A)\Bigr{)}\;+\;\frac{\rho_{2}}{2}\,\|\tilde{A}_{k}-A\|_{F}^{2}\Bigr{]}$
+$\quad+\;\alpha\,h(W_{k})\;+\;\frac{\rho_{1}}{2}\,\bigl{[}h(W_{k})\bigr{]}^{2},$
+where $\{\beta_{k}\}_{k=1}^{K}\in\mathbb{R}^{d\times d}$, $\{\gamma_{k}\}_{k=1}^{K}\in\mathbb{R}^{pd\times d}$ and $\alpha\in\mathbb{R}$ are estimates of the Lagrange multipliers; $\rho_{1}$ and $\rho_{2}$ are penalty coefficients, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
+Note that we set a single pair of $\lambda_{W},\lambda_{A}$ for all clients, as a generally effective pair can be found for datasets of the same dimension using FDBNL and prior studies [40, 41, 64].
+Due to the personalization problem, each $W_{k},A_{k}$ may have different dimensions or structural properties (e.g., connectivity).
+In this study, we consider only cases where all clients share identical dimensions and connectivity.
+However, extending the framework to handle different dimensions or varying connectivity structures is an interesting direction for future research, as discussed in § 6.
+We now define the iterative update rules of ADMM.
+Given initial values for $\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K}$, $W$, $A$, and the dual variables, the updates are as follows:
+For $W_{k},A_{k}$
+$\begin{split}(W_{k}^{(t+1)},A_{k}^{(t+1)})\;=\;\arg\min_{W_{k},A_{k}}\Bigl{[}\ell_{k}(W_{k},A_{k})&+\mu\|W_{k}-\tilde{W}_{k}^{(t)}\|^{2}+\mu\|A_{k}-\tilde{A}_{k}^{(t)}\|^{2}\\ &+\alpha^{(t)}\,h(W_{k})+\tfrac{\rho_{1}^{(t)}}{2}\,[h(W_{k})]^{2}+\lambda_{W}\|W_{k}\|_{1}+\lambda_{A}\|A_{k}\|_{1}\Bigr{]}.\end{split}$ (10)
+For $\tilde{W}_{k},\tilde{A}_{k}$:
+$\begin{split}(\tilde{W}_{k}^{(t+1)},\tilde{A}_{k}^{(t+1)})\;=\;\arg\min_{\tilde{W}_{k},\tilde{A}_{k}}\Bigl{[}&\mu\|W_{k}^{(t+1)}-\tilde{W}_{k}\|^{2}+\mu\|A_{k}^{(t+1)}-\tilde{A}_{k}\|^{2}+\mathrm{tr}\bigl{(}\beta_{k}^{(t)\top}(\tilde{W}_{k}-W^{(t)})\bigr{)}\\ &+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{W}_{k}-W^{(t)}\|^{2}+\mathrm{tr}\bigl{(}\gamma_{k}^{(t)\top}(\tilde{A}_{k}-A^{(t)})\bigr{)}+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{A}_{k}-A^{(t)}\|^{2}\Bigr{]}.\end{split}$ (11)
+For $W,A$:
+$\begin{split}(W^{(t+1)},A^{(t+1)})\;=\;&\arg\min_{W,A}\Bigl{[}\sum_{k=1}^{K}\Bigl{(}\mathrm{tr}\bigl{(}\beta_{k}^{(t)\top}(\tilde{W}_{k}^{(t+1)}-W)\bigr{)}+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{W}_{k}^{(t+1)}-W\|^{2}\Bigr{)}\\ &+\sum_{k=1}^{K}\Bigl{(}\mathrm{tr}\bigl{(}\gamma_{k}^{(t)\top}(\tilde{A}_{k}^{(t+1)}-A)\bigr{)}+\tfrac{\rho_{2}^{(t)}}{2}\|\tilde{A}_{k}^{(t+1)}-A\|^{2}\Bigr{)}\Bigr{]}.\end{split}$ (12)
+For Dual variables:
+$\begin{split}\beta_{k}^{(t+1)}&=\beta_{k}^{(t)}+\rho_{2}^{(t)}(\tilde{W}_{k}^{(t+1)}-W^{(t+1)}),\\ \gamma_{k}^{(t+1)}&=\gamma_{k}^{(t)}+\rho_{2}^{(t)}(\tilde{A}_{k}^{(t+1)}-A^{(t+1)}),\\ \alpha^{(t+1)}&=\alpha^{(t)}+\rho_{1}^{(t)}(\frac{1}{K}\sum_{k=1}^{K}h(W_{k}^{(t+1)})),,\\ \rho_{1}^{(t+1)}&=\phi_{1}\rho_{1}^{(t)},\\ \rho_{2}^{(t+1)}&=\phi_{2}\rho_{2}^{(t)},\end{split}$ (13)
+In the actual implementation of PFDBNL, we use a closed-form for $\tilde{W}_{k},\tilde{A}_{k}$ as derived below:
+$\tilde{W}_{k}^{(t+1)}\;=\;\frac{2\mu\,W_{k}^{(t+1)}\;+\;\rho_{2}^{(t)}\,W^{(t)}\;-\;\beta_{k}^{(t)}}{2\mu+\rho_{2}^{(t)}}.$
+$\tilde{A}_{k}^{(t+1)}\;=\;\frac{2\mu\,A_{k}^{(t+1)}\;+\;\rho_{2}^{(t)}\,A^{(t)}\;-\;\gamma_{k}^{(t)}}{2\mu+\rho_{2}^{(t)}}.$
+For the full procedure, we have summarized it in the supplementary material § 7.5.
+Similarly, in this work, we employ the L-BFGS method to solve the optimization problem.
+The procedure for PFDBNL is summarized in Algorithm 2.
+Algorithm 2 Personalized Federated Dynamic Bayesian Network Learning
+0:  Initial parameters $\rho_{1},\rho_{2},\alpha^{(1)},\beta_{1}^{(1)},\ldots,\beta_{K}^{(1)},\gamma_{1}^{(1)},\ldots,\gamma_{K}^{(1)}$; multiplicative factors $\phi_{1},\phi_{2}>1$; initial points $\{W_{k},A_{k},\tilde{W}_{k},\tilde{A}_{k}\}_{k=1}^{K},W,A$; participated clients at each round $j\leq K$
+1:  for $t=1,2,\ldots$ do
+2:     selected clients solve problem (10) first and solve (11) in parallel
+3:     Central server collects $\tilde{W}_{k_{1}}^{(t+1)},\ldots,\tilde{W}_{k_{j}}^{(t+1)},\tilde{A}_{k_{1}}^{(t+1)},\ldots,\tilde{A}_{k_{j}}^{(t+1)}$ from all clients
+4:     Central server solves problem (12)
+5:     Central server sends $W^{(t+1)}$ and $A^{(t+1)}$ to selected clients
+6:     Central server updates ADMM parameters $\alpha^{(t+1)},\rho_{1}^{(t+1)},\rho_{2}^{(t+1)}$ according to Eq.(13)
+7:     Each selected client updates $\beta_{k}^{(t+1)},\gamma_{k}^{(t+1)}$ according to Eq.(13)
+8:  end for
 4 Experiments
-Figure 1: An example result using 2Dbn for Gaussian noise data with $n=500$ samples, $d=5$ variables, an autoregressive order $p=3$, and $K=10$ clients.
+Figure 2: An example result using FDBNL for Gaussian noise data with $n=500$ samples, $d=5$ variables, an autoregressive order $p=3$, and $K=10$ clients.
+All clients have same $W,A$.
 We set the thresholds $\tau_{w}=\tau_{a}=0.3$.
 Our algorithm recovers weights close to the ground truth.
-In this section, we first study the performance of 2Dbn on simulated data generated by a linear SVAR structure [15].
-We then compare it against three linear baseline methods using three evaluation metrics, demonstrating the effectiveness of our proposed method.
-Figure 1 provides an illustrative example.
+In this section, we first evaluate the performance of FDBNL on simulated data generated from a linear SVAR structure [19].
+We then compare it against three linear baseline methods using three evaluation metrics to demonstrate the effectiveness of our proposed method.
+Figure 2 provides an illustrative example of homogeneous time series data.
+Next, we examine the performance of PFDBNL on heterogeneous time series data generated by a linear SVAR structure across different clients and also analyze its performance under partial client participation by uniformly sampling a subset of clients.
 Benchmark Methods.
-We compare our ADMM-based approach described in Sec.
-3, referred to as 2Dbn, with three other methods.
-The first is a baseline, denoted as Ave, which computes the average of the weighted adjacency matrices estimated by DYNOTEARS [36] from each client’s local dataset, followed by thresholding to determine the edges.
+We compare our federated approach, described in Sec.
+3.1 and referred to as FDBNL, with three other methods.
+The first baseline, denoted as Ave, computes the average of the weighted adjacency matrices estimated by DYNOTEARS [41] from each client’s local dataset, followed by thresholding to determine the edges.
 The second baseline, referred to as Best, selects the best graph from among those estimated by each client based on the lowest Structural Hamming Distance (SHD) to the ground truth.
-While this approach is unrealistic in practical scenarios (since it assumes knowledge of the ground truth), it serves as a useful point of reference.
-For additional context, we also consider DYNOTEARS applied to the combined dataset from all clients, denoted as Alldata.
-Note that the final graphs produced by Ave may contain cycles, and we do not apply any post-processing steps to remove them, as doing so could reduce performance.
-Since there is no official source code for DYNOTEARS, we re-implement it using only the numpy and scipy packages in about a hundred lines of code.
-This simpler, self-contained implementation eases understanding and reusability compared to available versions on GitHub.
+While this approach is impractical in real-world scenarios (since it assumes knowledge of the ground truth), it serves as a useful point of reference.
+For additional context, we also consider applying DYNOTEARS to the combined dataset from all clients, denoted as Alldata.
+Note that the final graphs produced by Ave may contain cycles for $W$, and we do not apply any post-processing steps to remove them, as doing so could degrade performance.
+Since there is no official source code for DYNOTEARS, we reimplement it using only the numpy and scipy packages in approximately a hundred lines of code.
+This simpler, self-contained implementation enhances both readability and reusability compared to existing versions on GitHub.
+For the personalized federated approach, referred to as PFDBNL, we compare its performance against FDBNL on heterogeneous time series data.
+Since no existing methods are specifically designed for learning heterogeneous DBNs, we focus on evaluating the improvements achieved by PFDBNL over FDBNL on a single example and demonstrate that PFDBNL produces reasonable results.
 Evaluation Metrics.
-We use three metrics to evaluate performance: Structural Hamming Distance, True Positive Rate (TPR), and False Discovery Rate (FDR).
-SHD measures the dissimilarity between the inferred graph and the ground truth, accounting for missing edges, extra edges, and incorrectly oriented edges [49].
-Lower SHD indicates closer alignment with the ground truth.
-TPR (also known as sensitivity or recall) quantifies the proportion of true edges correctly identified, calculated from true positives and false negatives [13].
-Higher TPR means the model identifies more true edges.
-FDR measures the proportion of false positives among all predicted edges, computed as the ratio of false positives to the sum of false positives and true positives [4].
+We use three metrics to evaluate performance: Structural Hamming Distance (SHD), True Positive Rate (TPR), and False Discovery Rate (FDR).
+SHD measures the dissimilarity between the inferred graph and the ground truth by accounting for missing edges, extra edges, and incorrectly oriented edges [57].
+A lower SHD indicates a closer alignment with the ground truth.
+TPR (also known as sensitivity or recall) quantifies the proportion of true edges correctly identified.
+It is computed as the ratio of true positives to the sum of true positives and false negatives [17].
+A higher TPR means the model successfully identifies more true edges.
+FDR measures the proportion of false positives among all predicted edges.
+It is defined as the ratio of false positives to the sum of false positives and true positives [5].
 A lower FDR indicates that most detected edges are correct.
-Together, these metrics provide a comprehensive understanding of the model’s accuracy in graph structure inference.
+Together, these metrics provide a comprehensive assessment of the model’s accuracy in graph structure inference.
+Furthermore, for heterogeneous time series data, since we have $K$ ground truth matrices $W_{k}$ and $A_{k_{i}}$, we evaluate performance separately for each client.
+Therefore, we report the mean SHD (mSHD), mean TPR (mTPR), and mean FDR (mFDR) across all clients.
+This is a reasonable evaluation metric, as it accounts for performance across all clients.
+4.1 Federated Dynamic Bayesian Network Learning
 Data Generation & Settings.
-We generate data following the Structural Equation Model described in Eq.
+For FDBNL, we generate data following the Structural Equation Model described in Eq.
 (2).
@@
 The intra-slice DAG is an Erdős-Rényi (ER) graph with a mean degree of 4, and the inter-slice DAG is an ER graph with a mean out-degree of 1.
-Although this results in very sparse graphs at high $d$, they remain connected under our settings [11].
-We set the base of the exponential decay of inter-slice weights to $\eta=1.5$.For the hyperparameters $\lambda_{w}$ and $\lambda_{a}$, we generate heatmaps to systematically identify their optimal values.
+Although this results in very sparse graphs at high $d$, they remain connected under our settings by Erdős and Rényi [13].
+We set the base of the exponential decay of inter-slice weights to $\eta=1.5$.
+For the hyperparameters $\lambda_{w}$ and $\lambda_{a}$, we generate heatmaps to systematically identify their optimal values.
 These values are documented in the Supplementary Materials (§7.3).
@@
 For DYNOTEARS (and thus Ave and Best), we follow the authors’ recommended hyperparameters.
-To avoid confusion, we summarize the data shape used in 2Dbn:
-  • For 2Dbn, $X$ and $Y$ are of size $K\times n^{\prime}\times d$ and $K\times n^{\prime}\times pd$, respectively, where $K$ is the number of clients, $n^{\prime}$ is the number of samples per client, $d$ is the dimensionality, and $p$ is the autoregressive order.
+To avoid confusion, we summarize the data shape used in FDBNL:
+  • For FDBNL, $X_{client}$ and $Y_{client}$ are of size $K\times n_{k}\times d$ and $K\times n_{k}\times pd$, respectively, where $K$ is the number of clients, $n_{k}$ is the number of samples for $k$-th client, $d$ is the dimensionality, and $p$ is the autoregressive order.
 Experiment Settings.
@@
 This setting evaluates adaptability and robustness as the data becomes more distributed.
-4.1 Varying Number of Variables
-In this section, we focus on DBNs with $n=5d$ samples evenly distributed across $K=10$ clients for $d=10,20$, and $n=6d$ for $d\in\{5,15\}$.
+4.1.1 Varying Number of Variables
+In this section, we focus on DBN with $n=5d$ samples evenly distributed across $K=10$ clients for $d=10,20$, and $n=6d$ for $d\in\{5,15\}$.
 Note that setting $n=5d$ or $6d$ ensures that each client has an integer number of samples.
@@
 Figure 3: Structure learning results for $W$ in a DBN with Gaussian noise for $d=5,10,15,20$ variables, an autoregressive order $p=1$, and $K=10$ clients.
+Each metric value indicates the mean performance across 10 different simulated datasets
 Figure 3 shows the results for the inferred $W$.
 The corresponding results for $A$ are provided in §10.
-Across all tested values of $d$, 2Dbn consistently achieves the lowest SHD compared to Ave and Best, and even outperforms Alldata for $d=5,15$.
+Across all tested values of $d$, FDBNL consistently achieves the lowest SHD compared to Ave and Best, and even outperforms Alldata for $d=5,15$.
 This advantage may be due to the need for more refined hyperparameter tuning in Alldata.
-In addition to a lower SHD, 2Dbn achieves a relatively high TPR, close to that of Alldata and higher than that of Best, indicating that 2Dbn accurately identifies most true edges.
-Moreover, 2Dbn attains a lower FDR than Ave and Best, emphasizing its superior reliability in edge identification under this challenging setup.
-4.2 Varying Number of Clients
+In addition to a lower SHD, FDBNL achieves a relatively high TPR, close to that of Alldata and higher than that of Best, indicating that FDBNL accurately identifies most true edges.
+Moreover, FDBNL attains a lower FDR than Ave and Best, emphasizing its superior reliability in edge identification under this challenging setup.
+4.1.2 Varying Number of Clients
 We now examine scenarios where a fixed total number of samples is distributed among varying numbers of clients.
@@
 For $\lambda_{w}$ and $\lambda_{a}$, we follow the recommendations of [41] for DYNOTEARS, Ave, and Best.
-For 2Dbn, we select the best regularization parameters from $[0.05,0.5]$ in increments of 0.05 by minimizing SHD.
+For FDBNL, we select the best regularization parameters from $[0.05,0.5]$ in increments of 0.05 by minimizing SHD.
 Figure 4: Structure learning for $W$ of a DBN with Gaussian noise for $d=10$ variables, $p=1$, and varying numbers of clients.
 There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
+Each metric value indicates the mean performance across 10 different simulated datasets
 Figure 5: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients.
 There are $n=512$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
-Figures 3 and 4 show the results for $W$ with $d=10$ and $d=20$.
+Each metric value indicates the mean performance across 10 different simulated datasets
+Figures 4 and 5 show the results for $W$ with $d=10$ and $d=20$.
 The results for $A$ are in §10.
 As $K$ increases, the TPRs of Ave and Best drop sharply, resulting in high SHDs.
-Although 2Dbn’s TPR also decreases as $K$ grows, it remains significantly higher than those of the other baselines.
-For example, with $d=20$ and $K=64$, 2Dbn achieves a TPR of 0.7, while Ave and Best achieve only 0.5 and 0.3, respectively.
-This highlights the importance of information exchange in the optimization process, enabling 2Dbn to learn a more accurate DBN structure even in highly distributed scenarios.
+Although FDBNL’s TPR also decreases as $K$ grows, it remains significantly higher than those of the other baselines.
+For example, with $d=20$ and $K=64$, FDBNL achieves a TPR of 0.7, while Ave and Best achieve only 0.5 and 0.3, respectively.
+This highlights the importance of information exchange in the optimization process, enabling FDBNL to learn a more accurate DBN structure even in highly distributed scenarios.
 Next, we consider the case of $n=256$ samples. The corresponding $W$-based results are shown below, and $A$-based results are in §10.
@@
 There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
+Each metric value indicates the mean performance across 10 different simulated datasets
 Figure 7: Structure learning for $W$ of a DBN with Gaussian noise for $d=20$ variables, $p=1$, and varying numbers of clients.
 There are $n=256$ total samples, distributed evenly across $K\in\{2,4,8,16,32,64\}$.
-From Figures 5 and 6, 2Dbn consistently attains the lowest SHD compared with Best and Ave.
-For $d=20$, the TPRs of Ave and Best continue to decline as $K$ increases, whereas 2Dbn’s TPR remains higher for all $K\geq 8$.
+Each metric value indicates the mean performance across 10 different simulated datasets
+From Figures 6 and 7, FDBNL consistently attains the lowest SHD compared with Best and Ave.
+For $d=20$, the TPRs of Ave and Best continue to decline as $K$ increases, whereas FDBNL’s TPR remains higher for all $K\geq 8$.
 At extremely high distributions, such as $K=64$ with $d=20$, each client has only 4 samples, leading to non-convergence for Ave and Best.
@@
 This suggests that the thresholding strategy used by Ave may lead to a high number of false positives.
-Nonetheless, 2Dbn’s TPR for $d=10$ remains close to that of Alldata as $K$ grows.
-In some cases, when $K$ or $d$ is small, the other baselines may outperform 2Dbn.
-For example, with $d=10,n=512$, Ave and Best achieve better SHDs than 2Dbn when $K\leq 16$.
+Nonetheless, FDBNL’s TPR for $d=10$ remains close to that of Alldata as $K$ grows.
+In some cases, when $K$ or $d$ is small, the other baselines may outperform FDBNL.
+For example, with $d=10,n=512$, Ave and Best achieve better SHDs than FDBNL when $K\leq 16$.
 This is expected since fewer clients and lower dimensionality provide each client with enough samples for accurate independent structure learning.
-However, as complexity grows (e.g., $d=20$), 2Dbn outperforms the baselines by maintaining a lower FDR and effectively leveraging distributed information.
+However, as complexity grows (e.g., $d=20$), FDBNL outperforms the baselines by maintaining a lower FDR and effectively leveraging distributed information.
 These findings highlight the importance of both method selection and hyperparameter tuning, as well as potential directions for future research in improving the efficiency of information exchange mechanisms for DBN learning.
+4.2 Personalized Federated Dynamic Bayesian Network Learning
+In this section, we first compare PFDBNL to FDBNL.
+For this example, we generate six pairs of $(W,A)$ for each client.
+Then, we simulate 10 datasets with $n_{k}=30$ for the $k$-th client.
+We compute the mean mSHD, mTPR, and mFDR after applying PFDBNL and FDBNL to each of these datasets.
+From Table 1, We observe that the mSHD of $W$ and $A$ in PFDBNL is half that of $W$ and $A$ in FDBNL.
+Furthermore, the mTPR of PFDBNL is twice as high as that of FDBNL.
+For mFDR, we observe that PFDBNL’s value is slightly lower than FDBNL’s, but the mFDR decreases by a factor of two.
+Hence, PFDBNL significantly improves performance across all evaluation metrics compared to FDBNL.
+To the best of our knowledge, no existing method is capable of learning Dynamic Bayesian Networks (DBNs) from heterogeneous time-series data.
+Given this gap, PFDBNL represents the first attempt to address this challenge and demonstrates its effectiveness in improving accuracy.
+Metric PFDBNL FDBNL
+mSHD $W$ 6.2 10.2
+$A$ 5.4 12.7
+mTPR $W$ 0.64 0.35
+$A$ 0.51 0.24
+mFDR $W$ 0.55 0.60
+$A$ 0.19 0.33
+Table 1: Comparison between FDBNL and PFDBNL with $d=5$, $K=6$, $n_{k}=30$ for $k\in K$.
+The expected degree of a node is 4 (both direction) with $p=1$.
+Each metric value indicates the mean performance across 10 different simulated datasets
+Data Generation & Settings.
+For PFDBNL, we generate data following the Structural Equation Model described in Eq.
+(2).
+This involves three steps: (1) Constructing weighted graphs $G_{W_{k}}$ and $G_{A_{k_{i}}}$ for each client.
+(2) Creating data matrices $X_{\text{client}}$ and $Y_{\text{client}}$ aligned with these graphs for each client.
+(3) Applying PFDBNL and evaluating its performance.
+Steps (1) and (2) follow the same data generation procedure as FDBNL, but partitioning of the dataset is not required.
+We use Gaussian noise with a standard deviation of 1 and set the base of the exponential decay of inter-slice weights to $\eta=1.5$.
+For the hyperparameters $\lambda_{W}$, $\lambda_{A}$, and $\mu$, we found that setting $\lambda_{W}=\lambda_{A}=0.1$ works well for $d=\{5,10\}$, while $\lambda_{W}=\lambda_{A}=0.01$ is effective for $d=\{15,20\}$ with $\mu=0.1$.
+We set $\phi_{1}=1.6$ and $\phi_{2}=1.1$, with initial values $\rho_{1}=\rho_{2}=1$.
+The initial Lagrange multipliers are set to zero.
+The data shape is the same as in the FDBNL experiment, but each client has a different $W_{k},A_{k_{i}}$.
+Experiment Settings.
+We consider the same two types of experimental settings as in the FDBNL experiment.
+First, we fix the number of clients $K$ and increase the number of variables $d$ while maintaining a consistent sample size $n$.
+This allows us to assess how well the methods scale with increasing dimensionality.
+Second, we fix the total sample size at $n=512$ and vary the number of clients $K$ from 2 to 32.
+This setting evaluates the adaptability and robustness of the methods as the data becomes more distributed.
+4.2.1 Varying Number of Variables
+In this section, we focus on learning of $W_{k}$ and $A_{k_{i}}$ with each $n_{k}=Kd$ samples for $d\in\{5,10,15,20\}$ across $K=6$ clients.
+Similarly to the example, we firstly generate $W_{k},A_{k_{1}}$ by using $p=1$ for all clients.
+Then we generate the dataset with $n_{k}$ sample sizes for $k$-th client.
+Furthermore, we exam our PFDBNL under two different connectivity level (one for low connectivity and one for high connectivity).
+We do not focused on mixed case in this study and leave it for future research direction as discussed in §6.
+For low connectivity, we set the expectation of degree of a node below half of number of variable as $\lfloor\frac{d}{2}\rfloor-\frac{d}{5}$.
+For high connective, we set expectation of degree of a node as $d-\frac{d}{5}$.
+Metric 5 10 15 20
+mSHD $W$ 4.4 16.8 44.7 79.2
+$A$ 2.8 10.9 29.6 32.2
+mTPR $W$ 0.61 0.49 0.55 0.53
+$A$ 0.81 0.53 0.31 0.31
+mFDR $W$ 0.72 0.67 0.51 0.53
+$A$ 0.35 0.47 0.23 0.34
+Table 2: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $d=5,10,15,20$ for low connectivity. Each metric value indicates the mean performance across 10 different simulated datasets
+Metric 5 10 15 20
+mSHD $W$ 6.2 13.6 52.0 110.0
+$A$ 5.4 8.3 41.0 31.0
+mTPR $W$ 0.64 0.64 0.71 0.76
+$A$ 0.51 0.52 0.31 0.27
+mFDR $W$ 0.55 0.54 0.44 0.59
+$A$ 0.19 0.21 0.36 0.19
+Table 3: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $d=5,10,15,20$ for high connectivity. Each metric value indicates the mean performance across 10 different simulated datasets
+From Tables 2 and 3, we observe that the mSHD for the low-connectivity case is much lower than in the high-connectivity case for $W$.
+However, the mSHD remains similar across all cases for $A$.
+For mTPR, we observe that performance in the high-connectivity scenario is significantly better than in the low-connectivity scenario.
+More importantly, while the mTPR for high-connectivity does not decrease with an increasing number of variables for $W$, it does decrease for $A$.
+Additionally, the mTPR for $A$ remains similar for $d\in\{10,15,20\}$, but in the low-connectivity case, it is 0.81, which is significantly higher than 0.51 in the high-connectivity case.
+This is not surprising, as identifying relationships in sparse networks is more challenging.
+For mFDR of $A$, we observe that it remains around 0.2 for $d\in\{5,10,20\}$ in the high-connectivity case.
+However, for the low-connectivity case, the mFDR increases to approximately 0.3.
+For $W$, the mFDR remains stable at around 0.54 in the high-connectivity scenario, which is significantly lower than in the low-connectivity case when $d\in\{5,10\}$.
+Overall, the mFDR in the high-connectivity case shows better results.
+We further explore the partial participation case under $d=10$.
+Since we only have six clients, we set the number of participating clients to $j=1,2,3,4,5$, while keeping all other settings unchanged.
+As shown in Table 4, the results remain stable regardless of the number of participating clients.
+The mSHD is approximately 13 to 14 for $W$ and around 8 for $A$.
+The mTPR remains about 0.60 for $W$ and 0.56 for $A$.
+The mFDR follows a similar pattern.
+Metric $j=1$ $j=2$ $j=3$ $j=4$ $j=5$
+mSHD $W$ 14.2 13.6 14.3 14.9 14.5
+$A$ 8.1 8.3 8.7 7.7 8.5
+mTPR $W$ 0.59 0.61 0.60 0.58 0.60
+$A$ 0.56 0.55 0.55 0.58 0.56
+mFDR $W$ 0.50 0.49 0.53 0.55 0.54
+$A$ 0.22 0.21 0.23 0.23 0.23
+Table 4: Partial participating case for $d=10$. Each metric value indicates the mean performance across 10 different simulated datasets
+4.2.2 Varying Number of Clients
+We now examine scenarios where a fixed total number of samples is distributed among varying numbers of clients.
+For $d\in\{10,20\}$, we generate $n=512$ samples in total.
+These samples are evenly distributed across $K\in\{2,4,8,16,32\}$ clients, meaning that each client receives $n_{k}=\{256,128,64,32,16\}$ samples, respectively.
+The mean degree of the intra-slice DAG is 4, counting edges in both directions.
+This may be a challenging setting because we need to learn different DBNs for 32 clients with only 16 samples per client.
+Metric $K=2$ $K=4$ $K=8$ $K=16$ $K=32$
+mSHD $W$ 11.3 20.6 17.4 18.2 19.4
+$A$ 6.7 10.5 10.7 10.8 14.8
+mTPR $W$ 0.68 0.44 0.51 0.51 0.49
+$A$ 0.63 0.52 0.48 0.48 0.36
+mFDR $W$ 0.49 0.62 0.54 0.54 0.59
+$A$ 0.17 0.31 0.23 0.28 0.37
+Table 5: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $K=2,4,8,16,32$ for $d=10$. Each metric value indicates the mean performance across 10 different simulated datasets
+Metric $K=2$ $K=4$ $K=8$ $K=16$ $K=32$
+mSHD $W$ 56.2 57.5 54.3 52.1 56.1
+$A$ 33.2 36.1 40.8 55.6 56.3
+mTPR $W$ 0.45 0.46 0.46 0.34 0.32
+$A$ 0.38 0.31 0.33 0.25 0.17
+mFDR $W$ 0.40 0.45 0.37 0.49 0.51
+$A$ 0.18 0.16 0.16 0.21 0.27
+Table 6: Comparison of metrics (mSHD, mTPR, mFDR) across dimensions $K=2,4,8,16,32$ for $d=20$. Each metric value indicates the mean performance across 10 different simulated datasets
+From Tables 5 and 6, we observe that PFDBNL consistently achieves mSHD values around 10–20 for $W$ and approximately 10 for $A$ when $d=10$.
+Notably, the mTPR of $W$ remains relatively stable as $K$ increases.
+For instance, at $K=32$, $W$ and $A$ obtain mTPRs of 0.49 and 0.36, respectively.
+This result is encouraging, even in scenarios where the number of clients ($K$) exceeds the number of samples per client.
+Meanwhile, the mFDR remains around 0.2–0.3 for $A$ but roughly 0.5 for $W$, suggesting that $A$ is more conservative, with fewer false discoveries.
+For $d=20$, although $W$ stabilizes at an mSHD of around 55, the performance of $A$ deteriorates from an mSHD of approximately 30 to 50 as $K$ increases, indicating greater sensitivity to high-dimensional settings.
+A similar pattern is observed in the mTPR: $W$ achieves approximately 0.45 when $K\in\{2,4,8\}$ but declines to about 0.3 for larger $K$.
+This drop is expected, as the sample size per client decreases while the number of parameters in the ground-truth matrix remains large.
+Overall, however, PFDBNL maintains stable and reasonable performance across different values of $K$, including the challenging scenario of $d=20,K=32,n_{k}=16$.
 5 Applications
@@
 We assume that each dataset corresponds to data collected from different hospitals or research centers, thereby motivating our federated approach.
-More information and data about DREAM4 are available at https://gnw.sourceforge.net/resources/DREAM4%20in%20silico%20challenge.pdf.
+More information and data about DREAM4 are available at https://gnw.sourceforge.net/resources/DREAM4%20in%20silico%20challenge.pdf
 Let $X^{g}_{t,r}$ denote the expression level of gene $g$ at time $t\in\{0,1,2,\ldots,20\}$ in replicate $r\in\{0,1,2,\ldots,R\}$.
@@
 [30].
-The federated setting of our 2Dbn approach includes $K=5$ clients, each with $R=2$ replicates.
+The federated setting of our FDBNL approach includes $K=5$ clients, each with $R=2$ replicates.
 Thus, each client contains a time-series dataset for 100 genes over 42 time steps.
 We found that small regularization parameters, $\lambda_{W}=\lambda_{A}=0.0025$, work well for all DREAM4 datasets.
+For our PFDBNL approach, we assume each dataset to be a client.
+Consequently, each client has 200 time steps with $d=100$ and $K=5$.
+We used regularization parameters $\lambda_{W}=\lambda_{A}=0.001$ and $\mu=0.1$.
 Lu et al.
-[26] evaluated various methods for learning these networks, including approaches based on Mutual Information(MI), Granger causality, dynamical systems, Decision Trees, Gaussian Processes (GPs), and Dynamic Bayesian Networks.
+[30] evaluated various methods for learning these networks, including approaches based on Mutual Information (MI), Granger causality, dynamical systems, Decision Trees, Gaussian Processes (GPs), and Dynamic Bayesian Networks (DBNs).
 Notably, we did not compare DYNOTEARS in this study, as its source code is not publicly available, and our implemented version did not achieve optimal results.
-The comparisons were made using AUPR (Area Under the Precision-Recall Curve) and AUROC (Area Under the Receiver Operating Characteristic Curve) across the 5 datasets.
-From Table.
-1, the GP-based method outperforms all others, achieving the highest mean AUPR (0.208) and AUROC (0.728).
-In terms of AUPR, our 2Dbn approach outperforms the TSNI (ODE-based) method in its non-federated version.
-Furthermore, the AUPR of 2Dbn is comparable to those of Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches.
-Importantly, the mean AUROC of 2Dbn surpasses those of GCCA (VAR-based), ARACNE (MI-based), and TSNI (ODE-based) methods, and is close to those of Ebdnet and VBSSMa (DBN-based).
-Overall, 2Dbn is comparable to other non-federated benchmarks.
+The comparisons were made using AUPR (Area Under the Precision-Recall Curve) and AUROC (Area Under the Receiver Operating Characteristic Curve) across the five datasets.
+From Table 7, the GP-based method outperforms all others, achieving the highest mean AUPR (0.208) and AUROC (0.728).
+In terms of AUPR, our FDBNL approach outperforms the TSNI (ODE-based) method in its non-federated version.
+Furthermore, the AUPR of FDBNL is comparable to those of Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches.
+Importantly, the mean AUROC of FDBNL surpasses those of GCCA (VAR-based), ARACNE (MI-based), and TSNI (ODE-based) methods, and is close to those of Ebdnet and VBSSMa (DBN-based).
+We also directly applied FDBNL, marked as FDBNL-d, to all datasets using the same settings as PFDBNL.
+From the table, we can see that if we directly use FDBNL, the results are not satisfactory.
+Therefore, we partitioned each dataset to ensure i.i.d.
+conditions for the use of FDBNL.
+We then tested PFDBNL, which performs well in heterogeneous situations.
+For PFDBNL, the AUPR is 0.034, which outperforms the TSNI (ODE-based) method in its non-personalized version.
+Additionally, the AUPR of FDBNL is only slightly greater than that of PFDBNL.
+Moreover, PFDBNL is also comparable to Ebdnet (DBN-based), GCCA (VAR-based), and ARACNE (MI-based) approaches.
+More importantly, the mean AUROC of PFDBNL surpasses that of ARACNE (MI-based) and is close to those of TSNI (ODE-based) and GCCA (VAR-based) methods.
+Overall, FDBNL and PFDBNL are comparable or even sometimes superior to other non-federated benchmarks.
 However, the GP-based method still achieves superior results in both AUPR and AUROC.
@@
 Unsurprisingly, the GP-based method outperforms all approaches because Gaussian Processes naturally handle nonlinear dynamics and are more adaptive to varying, complex relationships.
-In contrast, the 2Dbn model, although suitable for distributed data, might impose stronger parametric assumptions or face computational constraints that limit its ability to capture subtle nonlinear interactions.
+In contrast, the FDBNL and PFDBNL models, although well-suited for distributed data, may impose stronger parametric assumptions or face computational constraints that limit their ability to capture subtle nonlinear interactions.
 Thus, we believe GP-based methods could be extended to a distributed approach for structure learning, making them more applicable to real-world problems.
@@
 Method Type AUPR AUROC
-2Dbn FED-DBN 0.040 0.600
+FDBNL F-DBN 0.040 0.600
+PFDBNL PF-DBN 0.034 0.561
+FDBNL-d F-DBN 0.023 0.512
 Ebdnet [47] DBN 0.043 0.640
@@
 To conduct the experiments, we use simulated time series measurements corresponding to five different human subjects for each $d$ and compute the Average AUROC using the sklearn package.
-For the federated setting, we partition the 200 time steps among 5 clients ($K=5$).
+For the FDBNL setting, we partition the 200 time steps among five clients ($K=5$).
 Detailed information and descriptions of the data are available at https://www.fmrib.ox.ac.uk/datasets/netsim/index.html.
@@
 For $d=10$, we use the 2nd, 4th, 6th, 8th, and 10th subjects from Sim-2.mat.
-Finally, for $d=15$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-3.mat.
+For $d=15$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-3.mat.
+For the PFDBNL setting, we do not apply any partitioning; instead, we consider each subject as a client and we aaply our PFDBNL to 5 subjects.
+Thus, for $d\in\{5,10,15\}$, we have five clients, each containing 200 time steps.
+For $d=5$, we use the 2nd, 4th, 6th, 8th, and 10th subjects from Sim-1.mat.
+For $d=\{10,15\}$, we use the 1st, 3rd, 5th, 7th, and 9th subjects from Sim-2.mat and Sim-3.mat.
+To avoid any confusion, the difference between the two methods is as follows: (1) FDBNL is applied to each subject individually, where each subject is partitioned into five clients.
+As a result, we need to run five repetitions.
+(2) PFDBNL is directly applied to the five subjects without partitioning.
 Further details of this experiment are provided in the supplementary materials (see §7.6).
-We compare our method to the economy Statistical Recurrent Units (eSRU) proposed by Khanna and Tan [19] for inferring Granger causality, as well as existing methods based on a Multilayer Perceptron (MLP), a Long Short-Term Memory (LSTM) network [48], and a Convolutional Neural Network (CNN)-based model, the Temporal Causal Discovery Framework (TCDF) [34], on multivariate time series data for $d=15$, as examined by Khanna and Tan [19].
-As shown in Tab.2, our proposed 2DBNs achieve an AUROC of 0.74, outperforming the LSTM-based approach and approaching the performance of the CNN-based TCDF method.
-Even though 2DBNs do not surpass the MLP-based and eSRU-based methods, it is notable that the Federated version of our approach outperforms or closely matches several established non-Federated benchmarks.
-This outcome is reasonable, as the MLP and eSRU methods rely on deep architectures adept at modeling complex structural dependencies.
-Notably, our 2DBN method provides a new perspective on this problem by ensuring data security through its Federated approach.
-This capability is particularly important in scenarios involving sensitive or distributed datasets, as it allows for effective analysis without compromising privacy or data integrity.
-We have further elaborated on this aspect in Sec.6.
+We compare our method to the Economy Statistical Recurrent Units (eSRU) model proposed by Khanna and Tan [24] for inferring Granger causality, as well as existing methods based on a Multilayer Perceptron (MLP), a Long Short-Term Memory (LSTM) network [56], and a Convolutional Neural Network (CNN)-based model, the Temporal Causal Discovery Framework (TCDF) [39], on multivariate time series data for $d=15$, as examined by Khanna and Tan [24].
+As shown in Table 8, our proposed FDBNL achieves an AUROC of 0.74, outperforming the LSTM-based approach and approaching the performance of the CNN-based TCDF method.
+Although FDBNL does not surpass the MLP-based and eSRU-based methods, it is noteworthy that its federated version outperforms or closely matches several established non-federated benchmarks.
+Similarly, our proposed PFDBNL also achieves an AUROC of 0.74, demonstrating superior performance over the LSTM-based approach and approaching the CNN-based TCDF method.
+Notably, PFDBNL can be directly applied to multiple subjects, significantly reducing the computational cost associated with data preparation.
+While PFDBNL does not exceed the MLP-based and eSRU-based methods, its personalized version consistently outperforms or closely aligns with several established non-personalized benchmarks.
+This outcome is expected, as the MLP and eSRU methods rely on deep architectures that excel at modeling complex structural dependencies.
+Importantly, our FDBNL and PFDBNL methods offer a new perspective on this problem by ensuring data security through their federated and personalized approach.
+This capability is particularly crucial in scenarios involving sensitive or distributed datasets, as it enables effective analysis without compromising privacy or data integrity.
+We further elaborate on this aspect in Section 6.
 Method Averaged AUROC
@@
 eSRU [24] 0.84$\pm$0.03
-2DBNs 0.74$\pm$0.04
+FDBNL 0.74$\pm$0.04
+PFDBNL 0.74$\pm$0.05
 Table 8: Mean AUROC comparison of different methods for $d=15$.
 6 Discussion
-We proposed a federated framework for learning DBNs on horizontally partitioned data.
-Specifically, we designed a distributed DBN learning method using ADMM, where only model parameters are exchanged during optimization.
+We proposed a federated framework for learning Dynamic Bayesian Networks (DBNs) on homogeneous time series data and a personalized federated framework for heterogeneous time series data.
+Specifically, we designed a federated DBN learning method using ADMM, where only model parameters are exchanged during optimization.
 Our experimental results demonstrate that this approach outperforms other state-of-the-art methods, particularly in scenarios with numerous clients, each possessing a small sample size—a common situation in federated learning that motivates client collaboration.
+Additionally, we developed a personalized federated DBN learning method by incorporating a proximal operator and utilizing ADMM.
+Our experimental results indicate that the personalized approach improves accuracy while maintaining robust performance even with small sample sizes and a large number of clients.
+Furthermore, we demonstrate that partial client participation does not lead to a loss of accuracy.
 Below, we address some limitations of our approach and suggest potential directions for future research.
 Assumptions.
-We assumed that the structure of the DBN is fixed over time and is identical for all time series in the dataset (i.e., it is the same for all $m\in M$).
-Relaxing these assumptions could be useful in various ways, such as allowing the structure to change smoothly over time [44].
+We assume that the DBN structure, represented by $W,A$, remains constant over time and is the same for all time series in the dataset.
+Relaxing this assumption could be useful in various ways, such as allowing the structure to change smoothly over time [52].
 Another direction for future work is to investigate the behavior of the algorithm on non-stationary or cointegrated time series [32] or in scenarios with confounding variables [21].
 Federated Learning.
-The proposed ADMM-based approach to federated learning relies on a stateful setup, requiring each participating client to be involved in every round of communication and optimization.
+In our FDBNL, the ADMM-based approach to federated learning relies on a stateful setup, requiring each participating client to be involved in every round of communication and optimization.
+Even though we attempt to implement partial client participation for PFDBNL, applying the same strategy is meaningless since it only affects which clients join, without altering the fundamental process.
 This “always-on” requirement can be burdensome in real-world scenarios.
@@
 Ensuring that all such devices participate consistently and synchronously in every round is often impractical and can result in significant performance bottlenecks.
-An important future direction is to explore asynchronous techniques, which would enable stateless clients and facilitate cross-device learning.
-Furthermore, data may be vertically partitioned across clients, meaning that each client owns different variables but collectively aims to perform Bayesian Network Structure Learning (BNSL).
-Developing a federated approach tailored to this vertical setting represents another promising research direction.
+An important future direction is to apply other federated optimization techniques and explore asynchronous techniques, which would enable stateless clients and facilitate cross-device learning.
 Additionally, the ADMM procedure involves sharing model parameters with a central server, raising concerns about potential privacy risks.
@@
 To address this, exploring differential privacy techniques [12] to enhance the protection of shared model parameters is a critical avenue for future work.
-Moreover, our current work focuses on non-heterogeneous data.
-However, heterogeneous settings are more applicable to many real-world problems, as clients often differ in computational capabilities, communication bandwidth, and local data distributions.
+Personalized Learning.
+Heterogeneous settings are more applicable to many real-world problems, as clients often differ in computational capabilities, communication bandwidth, and local data distributions.
 These variations pose significant challenges for model convergence and performance.
-Several methods have been proposed to address these issues.
-For example, FedProx [24] incorporates a proximal term to stabilize and unify updates from clients with varying local training conditions, while FedNova [50] normalizes aggregated updates to mitigate the impact of heterogeneous local computational workloads.
+Thus, we propose the first methodology for learning Personalized DBNs by using the proximal operator.
+This framework can be easily extended to non-temporal cases, such as NOTEARS [64].
+As mentioned earlier, many potential personalized federated learning frameworks can be applied, such as meta-learning and clustered federated learning.
+Moreover, there are several interesting research directions to explore.
+In our paper, we study the personalized setting under a fixed $d$, i.e., the same number of variables for each client.
+A promising future direction is to address scenarios where $d$ varies across different clients.
+Similarly, for $p$, each client may have different time-lagged dependencies they wish to explore.
+Additionally, the case of mixed connectivity levels is an important research direction.
+For instance, in the presence of sparse networks, results for high-connectivity networks may be confounded, leading to missing edges.
+Several personalized federated learning methods have been proposed to address such issues.
+For example, FedNova [58] normalizes aggregated updates to mitigate the impact of heterogeneous local computational workloads.
+More importantly, asynchronous techniques are particularly important for heterogeneous cases.
+Even though our asynchronous technique does not lead to a loss of accuracy, applying asynchronous methods remains crucial and represents a promising research direction, especially in the context of reinforcement learning, as explored by Espeholt et al.
+[14].
 Nonlinear Dependencies.